### v1

# PROPS: Progressively Private Self-alignment of Large Language Models

Abstract
Alignment is a key step in developing Large Language Models (LLMs) using human feedback to ensure adherence to human values and societal norms. Dependence on human feedback raises privacy concerns about how much a labeler’s preferences may reveal about their personal values, beliefs, and personality traits. Existing approaches, such as Differentially Private SGD (DP-SGD), provide rigorous privacy guarantees by privatizing gradients during fine-tuning and alignment but can provide more privacy than necessary as human preferences are tied only to labels of (prompt, response) pairs and can degrade model utility. This work focuses on LLM alignment with preference-level privacy, which preserves the privacy of preference labels provided by humans. We propose PROPS (PROgressively Private Self-alignment), a multi-stage privacy preserving alignment framework where privately aligned models in previous stages can serve as labelers for supplementing training data in the subsequent stages of alignment. We present theoretical guarantees for PROPS as well as comprehensive validation using multiple models (Pythia and GPT) and datasets (AlpacaEval, Anthropic HH-RLHF, truthy-dpo-v0.1) to demonstrate the utility of PROPS over existing methods while still providing high privacy. For the same privacy budget, alignment via PROPS can achieve up to 3x higher win-rates compared to DP-SGD, and 2.5x higher win-rates compared to Randomized Response (RR) based alignment.
1 Introduction
The process of aligning LLMs relies on datasets comprising prompts, LLM-generated responses, and preference labels that indicate which response aligns better with human values, collectively referred to as preference data. The alignment approaches, including reinforcement learning with human feedback (RLHF) (Stiennon et al., 2020) and direct preference optimization (DPO) (Rafailov et al., 2024), leverage these preference datasets with ranked labels provided by human annotators.
Figure 1: (a) Randomized Response (RR) based alignment where human preferences in the dataset $\mathcal{D}$ are privatized using $R$ which are then used for alignment. (b) DP-SGD based alignment where differentially private gradients are used for model alignment. (c) Two stage PROPS framework: Dataset $\mathcal{D}$ is partitioned into disjoint subsets $(\mathcal{D}_{1},\mathcal{D}_{2})$. In Stage 11, preferences in $\mathcal{D}_{1}$ are privatized using RR, resulting in an intermediate aligned model $M_{1}$. In Stage 22, model $M_{1}$ is used to independently rank the responses in $\mathcal{D}_{2}$. We then obtain private labels for $\mathcal{D}_{2}$ which are derived from combining model’s predictions and RR via a maximum likelihood estimator (MLE). These progressively refined private preferences are then used for alignment to arrive at the final model $M_{2}$.
Motivation for Human Preference Privacy: Relying on human feedback for alignment brings up privacy concerns, particularly around the extent to which a labeler’s preferences might unintentionally expose their underlying values, identity, or personality. In medical applications, for example, publicly available case reports or anonymized symptoms can be used to prompt LLMs to generate diagnoses, with human feedback aligning these outputs to clinical best practices (Reddy, 2023; Han et al., 2024). Here, the preference data represents expert judgment about nuanced medical decisions and must remain private to protect the integrity and confidentiality of clinical expertise. In Wang et al. (2024), a large number of consultation records were used to construct a dataset to fine-tune an LLM to act as a chatbot physician, with GPT-4 used to retain records that highlighted professionalism, explainability, and emotional support. As the model was shown to be effective, this indicates that certain preferences of a doctor’s decision-making can be identified which could lead to potential exposure of a doctor’s preferences. Also, in policy analysis (Rao et al., 2023), publicly available survey questions or proposals can be used to elicit LLM-generated analyses, where policymakers’ feedback reveals sensitive interpretative insights that may require protection. Bakker et al. (2021), for example, observed that placing more emphasis on politics in their surveys to participants “resulted in self-reports of personality traits that were in some cases more aligned with preexisting political preferences.”
Recent work on privacy-preserving fine-tuning and alignment (see Section 2) typically treats prompts, responses, and human feedback as jointly private. However, in most alignment settings, only the human-provided labels or rankings are sensitive. While methods like differentially-private SGD (DP-SGD) protect the entire training tuple (as shown in Figure 1(b)), they hurt utility under stringent privacy requirements. This paper instead focuses on protecting human preference data. Figure 2 shows representative results that compare the method proposed in this work with conventional privacy preserving techniques including DP-SGD and Randomized Response (RR); the figure shows that for the same privacy guarantee, models aligned by our method provide higher quality responses compared to DP-SGD and RR based alignment.
Main Contributions: Motivated by the above observations, we study the problem of aligning LLMs with preference privacy. Specifically, we investigate two notions of privacy: a) preference-level privacy, and b) labeler-level privacy. Preference level privacy ensures that for any tuple $(x,y_{1},y_{2})$, where $x$ denotes the prompt and $y_{1},y_{2}$ denote the LLM-generated responses, the individual human-preference $\ell^{*}$ (which denotes whether $y_{1}$ or $y_{2}$ is preferable) does not significantly impact the aligned model. Formally, we leverage the existing notion of Differential Privacy (DP) Dwork et al. (2014), and use it to formalize the notion of $(\epsilon,\delta)$-preference-level DP, where $(\epsilon,\delta)$ represent the privacy budgets. The notion of labeler-level privacy (also commonly referred to as user-level privacy in the DP literature) hides the presence/absence of any individual human labeler and protects all the labels annotated by the labeler. We summarize and highlight the main contributions and novel aspects of this work:
  • PROPS for Private Alignment. We introduce Progressively Private Self-Alignment (PROPS), a multi-stage algorithm for alignment that improves privacy and utility. Instead of processing the entire perturbed dataset at once, PROPS divides alignment into stages. In the $k^{\text{th}}$ stage, for $(k>1)$, the model from the previous stage ($M_{k-1}$) is used along with non-private prompts, responses from a new batch, and noisy labels perturbed using Randomized Response (RR) $\ell_{\text{RR}}$. $M_{k-1}$ generates its own rankings $\ell_{M_{k-1}}$ which, combined with $\ell_{\text{RR}}$, are used to compute maximum-likelihood estimates (MLEs) for alignment. This staged process leverages intermediate models to improve preference labeling and reduce reliance on noisy labels, enhancing alignment quality while maintaining privacy. As the model evolves over $K$ stages, alignment becomes progressively refined. PROPS effectively balances privacy preservation with performance, offering a novel alignment framework.
  • Theoretical Insights. We study the utility-privacy tradeoffs for PROPS by analyzing the Sub-optimality gap Chowdhury et al. (2024a) defined as the difference between between the weights of a non-privately trained model with the weights of a privately trained model with a privacy budget $\epsilon$ in Section 3.1 (Theorem 1). The upper bound on this gap shows that PROPS (which uses MLE combining in the second stage) is no worse than vanilla RR and performs better as long as the intermediate model gets better at predicting preferences.
  • Empirical Evaluation. We conducted a comprehensive set of experiments to evaluate the impact of preference-level differential privacy (DP) on DPO-based alignment across various privacy settings and models (Pythia-1B, GPT2-Large, and GPT2-Medium). Our results show that in the high privacy regime $(\epsilon=0.1)$, our method, PROPS, achieves up to 2.5x preference gain for PROPS vs RR in win-tie-loss rates and up to 3x win-tie-loss rate preference gain for PROPS vs DP-SGD based alignment on truthy-dpo-v0.1, HH-RLHF and AlpacaEval datasets. We refer the readers to Section 4 and Section A.6 for detailed experimental results.
Figure 2: (a) Win-Tie rate evaluation of PROPS vs RR and PROPS vs DP-SGD-aligned models on the truthy-dpo-v0.1 dataset for GPT2-Large and GPT2-Medium models, demonstrating the advantages of preference-level privacy with PROPS, particularly in high-privacy regimes. (b) Prompt-Response pairs generated by GPT2-Large model with PROPS, DP-SGD and $R$-based alignment for different privacy regimes.
Related works & Limitations: We next provide an overview of LLM alignment and the associated privacy risks in using human preference data. We also discuss related works on this problem and the limitations of existing methods for privacy-preserving alignment.
LLM Alignment: Training LLMs typically involve three key stages: pre-training, supervised fine-tuning, and alignment. Among these, alignment is particularly important as it guides LLMs to produce responses that align with societal norms and human preferences. The alignment process relies on a dataset $\mathcal{D}$ consisting of $n$ samples, each containing a prompt $x$, LLM-generated responses $(y_{1},y_{2})$, and a human-preferred label $\ell^{*}$, collectively referred to as preference data. Two conventional methods for alignment, Reinforcement Learning with Human Feedback (RLHF) (Stiennon et al., 2020) and Direct Preference Optimization (DPO) (Rafailov et al., 2024), utilize preference datasets with ranked labels provided by human annotators. While these aligned models improve the quality of generated responses, they introduce privacy risks, particularly concerning the identity of the annotators and their associated preferences. Recent work shows membership inference attack on preference data for LLM alignment (Feng et al., 2024), and highlights the vulnerabilities of using human annotated preference data during alignment. To mitigate the privacy risks associated with human-annotated preference data, the notion of differential privacy (DP) (Dwork et al., 2014) has recently been explored for fine-tuning and alignment of LLMs. For example, Yu et al. (2021) applied DP to fine-tuning by introducing privacy guarantees for smaller, appended parameters such as LoRA and adapters. Singh et al. (2024) introduced a two-stage fine-tuning process, and Yu et al. (2024) addressed the privacy-preserving alignment challenge by ensuring DP protection for users’ prompts against labelers during the generation of preference datasets for alignment. Additionally, Wu et al. (2023) proposed applying DP to RLHF by splitting the dataset into three disjoint sets to ensure DP at each stage of RLHF. Moreover, Feng et al. (2024) investigated the vulnerability of LLMs aligned using human preference datasets to membership inference attacks (MIAs), and provided empirical evidence that DPO models are more vulnerable to MIA compared to RLHF based models.
Private Alignment: To address privacy threats, the idea of Differential Privacy (DP) (Dwork et al., 2014) has been integrated into the LLM training process, including pre-training, fine-tuning, and alignment, to protect the privacy of preference data. To achieve DP, current state-of-the-art approaches utilize DP-SGD (Abadi et al., 2016), which privatizes gradients to ensure the privacy of the entire preference data throughout the training process. Previous literature has explored key aspects relevant to this study: (Chowdhury et al., 2024b) examined privacy preserving reward estimation for RLHF methods, while (Chowdhury et al., 2024a) investigated the robustness of DPO with noisy preference datasets, providing the foundational basis for this study. However, to the best of our knowledge, the concept of Label Differential Privacy (Label-DP) (Ghazi et al., 2021b) has not been explored in the context of alignment. Label-DP could enable models to achieve high utility while reducing privacy leakage, particularly when protecting the labels generated by human annotators. While DP-SGD offers privacy guarantees for the entire dataset, including the prompt $x$, LLM-generated responses $(y_{1},y_{2})$, and human-generated label $\ell^{*}$; we notice that it is primarily the labels in the preference data that pose privacy risks concerning the human labelers. Based on this critical observation, we propose a novel framework for progressively private LLM alignment that ensures a certain level of privacy for human labelers. Our framework’s primary goal is to protect human-generated labels in preference datasets while achieving a better privacy-utility trade-off. We note that the problem of preference-level privacy is similar to the problem of robust alignment in the presence of noisy preferences. Specifically, recent works, including Mitchell (2023), Chowdhury et al. (2024a) and Chowdhury et al. (2024b) study the robustness of alignment when the human-annotated labels are intrinsically noisy. The distinction, however, is the following: in our setting, the injected noise (and more importantly, the parameters of the privacy-preserving mechanisms (detailed in the next Section)) are known and can be controlled as a function of the privacy parameters.
2 Preliminaries on Alignment & Privacy
We start with a preference dataset $\mathcal{D}$ with $n$ samples where the $i^{th}$ sample can be expressed as $(x_{i},y_{1}^{i},y_{2}^{i},\ell_{i}^{*})$ where $x_{i}$ is the prompt, $y_{1}^{i},y_{2}^{i}$ are two LLM generated responses and $\ell_{i}^{*}$ is the human chosen label that can be defined as:
$\ell_{i}^{*}=\begin{cases}1,\quad\text{if }y_{1}^{i}\text{ is preferred over the response }y_{2}^{i}\\ 0,\quad\text{otherwise}.\end{cases}$
For the ease of notation, we define $y_{p}$ as the preferred response and $y_{np}$ as the not-preferred response, and suppress the index $i$. Specifically, if $\ell^{*}=1$, then we will have $y_{p}=y_{1},y_{np}=y_{2}$, and for the case $\ell^{*}=0,y_{p}=y_{2},y_{np}=y_{1}$. For a prompt $x$ in the dataset $\mathcal{D}$ with $y_{p}$ response preferred over the response $y_{np}$, we define the DPO (instance specific) loss as:
$\textsf{loss}^{*}(x,y_{p}\succ y_{np})$ $=\log\sigma\bigg{\{}\beta\log\frac{\pi_{\theta}(y_{p}|x)}{\pi_{\text{ref}}(y_{p}|x)}-\beta\log\frac{\pi_{\theta}(y_{np}|x)}{\pi_{\text{ref}}(y_{np}|x)}\bigg{\}}$
$=\mathds{1}(\ell^{*}=1)\textsf{loss}(x,y_{1}\succ y_{2})\mathrel{+}\mathds{1}(\ell^{*}=0)\textsf{loss}(x,y_{2}\succ y_{1}),$ (1)
where $\pi_{\theta}$ and $\pi_{\text{ref}}$ represent the current version of the LLM being optimized and the initial version of the LLM prior to alignment respectively, and $\beta$ is a constant used to control the penalty for how much $\pi_{\theta}$ diverges from $\pi_{\text{ref}}$. The instant-specific true loss mentioned in Equation equation 1 represents the loss for every prompt $x$ in preference data $\mathcal{D}$, therefore the expected DPO loss can be defined as:
$\mathbb{E}[\textsf{loss}(x,y_{1},y_{2},\ell^{*})]$ $=\mathbb{E}_{(x,y_{1},y_{2},\ell^{*})\sim\mathcal{D}}\big{\{}\mathds{1}(\ell^{*}=1)\cdot\textsf{loss}(x,y_{1}\succ y_{2})$
$+\mathds{1}(\ell^{*}=0)\cdot\textsf{loss}(x,y_{2}\succ y_{1})\big{\}}.$ (2)
2.1 Privacy for Alignment
The notion of differential privacy (DP) Dwork et al. (2014) has been adopted in the alignment frameworks to ensure that the presence or absence of a single sample in a preference dataset does not significantly alter the outcome of the model.
Definition 1 ($(\epsilon,\delta)$ Differential Privacy).
For all pairs of neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ that differ by a single entry, i.e., $\mathcal{D}-\mathcal{D}^{\prime}||_{1}\leq 1$, a randomized algorithm $\mathcal{M}$ with an input domain of D and output range $\mathcal{R}$ is considered to be $(\epsilon,\delta)$-differentially private, if $\forall\mathcal{S}\subseteq\mathcal{R}$:
$\small\mathbb{P}[\mathcal{M}(\mathcal{D})\in\mathcal{S}]\leq e^{\epsilon}\cdot\mathbb{P}[\mathcal{M}(\mathcal{D}^{\prime})\in\mathcal{S}]+\delta.$
We next introduce the notion of preference-level privacy and explain how it can be expanded to labeler-level privacy. Specifically, preference level privacy ensures that the LLM after alignment should not be significantly impacted by a change in a single preference.
Definition 2 ($(\epsilon,\delta)$-Preference level DP).
For all neighboring datasets $\mathcal{D}$ and $\mathcal{D^{\prime}}$ that differ by one preference ranking (i.e. $\{x_{i},y_{1}^{i},y_{2}^{i},\ell_{i}\}\in\mathcal{D}$ and $\{x_{i},y_{1}^{i},y_{2}^{i},(1-\ell_{i})\}\in\mathcal{D^{\prime}}$, a model after performing an alignment procedure $M$, whose output domain $S$ consists of all possibly aligned models, will satisfy $(\epsilon,\delta)$-preference level DP if :
$\mathbb{P}[M(D)\in S]\leq e^{\epsilon}\cdot\mathbb{P}[M(D^{\prime})\in S]+\delta.$ (3)
From Preference-level DP to Labeler-level DP: Preference-privacy protects individual labeling actions, such as rating a single prompt-response pair. However, when a labeler annotates multiple prompt-response pairs across the dataset, labeler-privacy guarantees become essential. This distinction between preference-privacy and labeler-privacy has been well recognized in the literature (McMahan et al., 2017; Liu et al., 2020; Levy et al., 2021). To extend preference-privacy guarantees to the labeler-privacy, privacy accounting and composition techniques can be adopted. For instance, if a labeler contributes to $k$ labeled examples in the dataset $\mathcal{D}$ with $(\epsilon,0)$-preference privacy, the Basic Composition theorem (Dwork et al., 2014) implies a labeler-privacy guarantee of $(k\epsilon,0)$. To limit cumulative privacy loss, it is important to operate in high-privacy regimes (i.e., with small privacy budgets). Notably, we observe that PROPS performs significantly better in such regimes, where the resulting labeler-privacy guarantees are stronger. Composition is the key technique that we use to obtain labeler-level priacy from preference-level privacy when individual labeler labels more than a single prompt-response pair. For instance, in dataset $\mathcal{D}$, if any labeler labels $k$ prompt-response pairs, with Basic Composition, $(\epsilon,0)$-Preference DP will satisfy $(k\epsilon,0)$-Labeler DP. However, more sophisticated methods, such as Advanced Composition (Dwork et al., 2014), Adaptive Composition Rogers et al. (2016), and the Moments Accountant Abadi et al. (2016) can be used to obtain tighter bounds depending on the application’s privacy requirements. With a small failure probability of $\delta^{\prime}$, Advanced composition provides ($\epsilon_{\text{Labeler}},\delta_{\text{Labeler}}$) Labeler DP where $\epsilon_{\text{Labeler}}=k_{i}\epsilon^{2}+\epsilon\sqrt{2k_{i}\log(\frac{1}{\delta^{\prime}})},$ and $\delta_{\text{Labeler}}=\delta^{\prime}$.
2.2 Limitations of Existing Approaches for Differentially Private Alignment
The two primary approaches for achieving private alignment in machine learning are Randomized Response (RR) and Differentially Private Stochastic Gradient Descent (DP-SGD). While Randomized Response (RR) is a simple baseline for achieving $\epsilon$-preference-level differential privacy (DP), RR perturbs the preference labels in the dataset $\mathcal{D}=\{x,y_{1},y_{2},\ell^{*}\}$. The perturbed dataset output by the RR mechanism is $\{x,y_{1},y_{2},\ell_{RR}\}$, where the label $\ell_{RR}$ is flipped with probability $\gamma_{\epsilon}=\dfrac{1}{1+e^{\epsilon}}$:
$\small\ell_{RR}=\begin{cases}\ell^{*},\text{ with probability }(1-\gamma_{\epsilon})=\dfrac{e^{\epsilon}}{1+e^{\epsilon}}\\ 1-\ell^{*},\text{ with probability }\gamma_{\epsilon}=\dfrac{1}{1+e^{\epsilon}}.\end{cases}$
Though RR is simple to implement and ensures strong privacy guarantees, it introduces significant noise to the labels, which can degrade alignment quality, especially in small datasets or high-privacy regimes. Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al., 2016) ensures $(\epsilon,\delta)$-DP by gradient perturbation during alignment. In each round, the gradients $(\mathbf{\bar{g}})$ are clipped with a clipping threshold $(C)$ and perturbed with Gaussian Noise $(\mathcal{N}(0,{\sigma}^{2}{C}^{2}I))$ where, noise scale $\sigma=\sqrt{2\log(1.25/\delta)}/\epsilon$. In scenarios requiring stringent privacy constraints such as alignment, DP-SGD often reduces model utility, since it perturbs the gradient privatizing the prompt-response pairs as well as the human annotated labels. This tradeoff limits its effectiveness of alignment. These limitations of aforementioned methods highlight the need for more sophisticated approaches that achieve a better balance between privacy preservation and model performance. To address this, we propose Progressively Private Self-Alignment (PROPS), a novel framework that leverages intermediate alignment stages to improve utility while ensuring preference-level privacy. For more related works on LLM alignment and Privacy, we refer the readers to the Section A.1.
3 PROPS: Progressively Private Self-Alignment
Figure 3: Key building blocks of PROPS framework. The figure illustrates the label generation of PROPS: In the first round, the human annotated labels $\ell^{*}$ are perturbed using $R$ ($\ell_{\text{RR}}$) which are then used to align model $M_{1}$. In every $(k+1)^{\text{th}}$ round, model $M_{k}$ predicted labels ($\ell_{M_{k}}$) and $R$-based labels $\ell_{\text{RR}}$ are then selected based on MLE to achieve labels $\ell_{\text{PROPS}}$.
In this Section, we present Progressively Private Self-Alignment algorithm (PROPS), which is the main technical contribution of this paper. Our key idea is to preserve the privacy of the labeler generated preferences during the multi-stage alignment process for better privacy-utility trade-off. To facilitate understanding, we first describe PROPS in a two-stage $(K=2)$ setting and the generalization to arbitrary number of stages is straightforward. We begin with the preference dataset $\mathcal{D}$ consists of $n$ samples. Each sample is represented as $(x,y_{1},y_{2},\ell^{*})$, where $x$ is the prompt, $(y_{1},y_{2})$ are the large language model (LLM) generated responses, and $\ell^{*}$ is the human labeler’s preference. We partition this dataset into two halves, denoted as $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. Let us perturb the labels of each entry using the RR mechanism, i.e., the labels are flipped with probability $\gamma_{\epsilon}=1/(1+e^{\epsilon})$.
Stage $1$: In the first stage, we use the dataset $D_{1}$ (with perturbed labels using RR) and use it to align a fine-tuned model via DPO. Let’s denote the resulting model as $M_{1}$. First note that since the training was done on private (perturbed preferences), due to post-processing the model $M_{1}$ can be used in subsequent stages without additional leakage.
Stage $2$: In this stage, we use the dataset $D_{2}$, and model $M_{1}$ (of the previous stage) to label/rank the preference of each prompt/response-pairs. Note that this procedure only requires the prompt and response pairs (and not the ground-truth human preferences); thus, this does not cause any additional privacy leakage.
Lemma 1.
For all $\delta^{\prime}\geq 0$, PROPS framework satisfies $(\epsilon,0)$-Preference DP. If no labeler labels more than k prompt-response pairs in dataset $\mathcal{D}$, then PROPS satisfies $(\epsilon_{\text{Labeler}},\delta_{\text{Labeler}})$ Labeler DP, where $\epsilon_{\text{Labeler}}=k\epsilon^{2}+\epsilon\sqrt{2k\log(\frac{1}{\delta^{\prime}})},\delta_{\text{Labeler}}=\delta^{\prime}$.
Let us denote the corresponding label obtained from the model $M_{1}$ for a prompt as $\ell_{M_{1}}$. To summarize, at this point, for each tuple $(x,y_{1},y_{2})$, we have access to $(\ell_{RR},\ell_{M_{1}})$. Viewing these as two noisy observations of the ground truth $\ell^{*}$, it is natural to ask the following: can one combine these predictions to obtain a better prediction about the ground truth preference?
Note that we know the error rate of RR ($\gamma_{\epsilon}$), however, we don’t know the error rate of the model $M_{1}$ (say $\gamma_{M_{1}}$, which denotes the probability with which the model $M_{1}$ makes errors). However, if we knew the error rate of the model $M_{1}$ (or an estimate for $\gamma_{M_{1}}$), then we could then use a combining approach (e.g., the maximum-likelihood estimator or MLE) to design a potentially better estimate of the ground truth label for alignment. In fact, it is not too difficult to work out the MLE combiner using two noisy observations. Assuming that the RR noise and the noise induced by the model $M_{1}$ are independent, the MLE statistic (log-likelihood ratio) can be written and simplified as follows:
$\small\Lambda(\ell_{RR},\ell_{M_{1}})$ $=\log\left(\frac{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=0)}{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=1)}\right)$
$=(-1)^{\ell_{RR}}\cdot\log\left(\frac{1-\gamma_{\epsilon}}{\gamma_{\epsilon}}\right)+(-1)^{\ell_{M_{1}}}\cdot\log\left(\frac{1-\gamma_{M_{1}}}{\gamma_{M_{1}}}\right).$ (4)
The above then yields the methodology one can use for creating a new label for each prompt as:
$\ell_{{\text{PROPS}}}(\ell_{RR},\,\ell_{M_{1}})=\begin{cases}1,\quad\text{if }\Lambda(\ell_{RR},\,\ell_{M_{1}})\leq 0\\ 0,\quad\text{if }\Lambda(\ell_{RR},\,\ell_{M_{1}})>0.\end{cases}$ (5)
With access to the new label estimates, $\ell_{\text{PROPS}}$ (for all samples in the set $\mathcal{D}_{2}$), we then train $M_{1}$ using DPO to obtain a new model $M_{2}$. This procedure can be repeated in a multi-stage setting by replacing $M_{1}$ by $M_{k-1}$ which is then trained on PROPS labels to obtain the model $M_{k}$ for the next stage.
Estimating $\gamma_{M_{1}}$: We next present an interesting approach to estimate $\gamma_{M_{1}}$, the rate at which the model $M_{1}$ predicts incorrect labels. Under the assumption that the model $M_{1}$ independently flips the ground truth labels with probability $\gamma_{M_{1}}$, then we can write the output of the model and RR mechanisms respectively as follows:
$\ell_{RR}=\ell^{*}\oplus U,\quad\ell_{M_{1}}=\ell^{*}\oplus V,$ (6)
where $U\sim\text{Bern}(\gamma_{\epsilon})$ and $V\sim\text{Bern}(\gamma_{M_{1}})$. Thus, estimating $\gamma_{M_{1}}$ is equivalent to estimating the parameter of the Bernoulli random variable $V$. Note that we have $|\mathcal{D}_{2}|$ observations (one for each sample in the second half of the dataset). If we compute
$\small\mu_{M_{1}}$ $=\frac{\sum_{i}\ell^{(i)}_{M_{1}}\oplus\ell^{(i)}_{RR}}{|\mathcal{D}_{2}|},$ (7)
which represents the number of disagreements between the labels predicted by RR and the model $M_{1}$, this value in fact converges to the the expected value of $\mathbb{E}[U\oplus V]=\gamma_{M_{1}}(1-\gamma_{\epsilon})+\gamma_{\epsilon}(1-\gamma_{M_{1}}))$. Since we know $\gamma_{\epsilon}$, we can then use it to compute the unknown parameter $\gamma_{M_{1}}$. This leads us to propose the following estimator for $\gamma_{M_{1}}$:
$\small\text{(Estimate of $\gamma_{M_{1}}$)}~~~\hat{\gamma}_{M_{1}}$ $=\frac{\mu_{M_{1}}-\gamma_{\epsilon}}{1-2\gamma_{\epsilon}}.$ (8)
The detailed proof of (8) and the fact that above estimator is unbiased is presented in Section A.3. In addition, in Section A.3, we provide experimental evidence to verify the efficacy of our proposed approach and the assumption about $M_{1}$, by comparing our proposed estimation procedure with the scenario where an oracle provides access to the true labels. Our results indicate that our procedure is valid, as the gap between the estimated and “ground-truth” $\gamma_{M}$ is consistently small across various privacy budgets.
3.1 Theoretical Results for PROPS
In this section, we analyze the Sub-optimality gap, which captures the gap between the optimal non-private DPO policy parameters $\theta^{*}$, and the policy parameters obtained through two-stage PROPS $\hat{\theta}_{\text{PROPS}}$. To obtain a bound on the Sub-optimality gap, we follow similar assumptions as those made in Chowdhury et al. (2024a), including expressing the model as a log-linear policy and assuming smoothness by placing bounds on the policy and its gradients. This gap is formalized in Theorem 1; we provide more details on the assumptions and derivation in Section A.5.
Theorem 1.
Under the smoothness assumption described above, for a log-linear policy class, 2-stage PROPS achieves a sub-optimality gap bounded as:
$\small\underbrace{\left\lVert\hat{\theta}_{\text{PROPS}}-\theta^{*}\right\rVert}_{\text{Sub-Optimality Gap}}\leq\mathcal{O}\left(\frac{\sqrt{\kappa}}{\gamma\beta(1-2\cdot\min(\gamma_{M_{1}},\gamma_{\epsilon}))}\sqrt{\frac{d}{n_{2}}}\right),$
at least with probability of $(1-\delta)$ where, $\delta\in(0,1]$, $\kappa$ is a constant, $n_{2}$ is the number of samples in the second-stage, $d$ denotes the dimensionality of the feature space and $\kappa$ represents the relative feature coverage between $\pi_{\theta}$ and $\pi_{\text{ref}}$ (i.e. fine-tuned policy).
The upper bound shows that PROPS (which uses MLE combining in the second stage) is always better than vanilla RR as long as $\gamma_{M_{1}}<\gamma_{\epsilon}$. It also indicates how the amount of training data in a particular stage of PROPS can affect performance. If more data is used for second-stage training (larger $n_{2}$) compared to that of the first stage, the sub-optimality gap decreases, which may not lead to a sufficiently aligned initial model $M_{1}$. Conversely, allocating a larger portion of data for first-stage training (larger $n_{1}$) may yield a stronger initial model $M_{1}$ but reduces the size of $n_{2}$, potentially increasing the sub-optimality gap. To strike a balance between these trade-offs, we split the full dataset in half for both stages (i.e. $n_{1}=n_{2}=n/2$ for a dataset $\mathcal{D}$ of $n$ preference samples). This helps ensure $M_{1}$ is sufficiently aligned while maintaining reliable performance in the second-stage.
PROPS Algorithm & Remarks: We present the main algorithm of this paper (PROgressively Private Self-alignment) PROPS in Algorithm 1 and present a set of remarks regarding this algorithm:
Algorithm 1 PROPS: PROgressively Private Self-alignment
Inputs: Fine-tuned Model $M_{0}$, Dataset $\mathcal{D}$, Stages $K$, Privacy parameters $(\epsilon,\delta)$
Output: Aligned model $M_{K}$
 Perform RR on $\mathcal{D}$, such that $\mathcal{D}\xrightarrow{RR(\gamma_{\epsilon})}\mathcal{D}^{\prime}$
 Partition the dataset $\mathcal{D}^{\prime}$ into $K$-disjoint datasets such that $\mathcal{D}^{\prime}=D_{1}\cup D_{2}\cup\ldots\cup D_{K}$
 Flip labels using Randomized Response (RR) with probability $\gamma_{\epsilon}=\frac{1}{1+e^{\epsilon}}$
 Align model $M_{0}$ with dataset $D_{1}$ to obtain $M_{1}$
for $k=2,3,\ldots,K$ do
  Generate labels $\ell_{M_{k-1}}^{D_{k}}$ for dataset $D_{k}$ using the model $M_{k-1}$ (from previous stage)
  Obtain $\ell_{RR}^{D_{k}}$ and $\ell_{M_{k-1}}^{D_{k}}$ and obtain the maximum likelihood estimator (MLE) $\Lambda$ according to equation 9 and generate label as:
$\small\ell_{\text{PROPS}}^{D_{k}}=\begin{cases}1,\quad\text{ If }\Lambda\leq 0\\ 0,\quad\text{ If }\Lambda>0.\end{cases}$
  Align model $M_{k-1}$ on dataset $D_{k}$ with PROPS labels ($\ell_{\text{PROPS}}^{D_{k}}$) to get model $M_{k}$
end for
Output: Aligned model $M_{K}$ on dataset $\mathcal{D}$
Remark 1 PROPS for RLHF. While we have presented PROPS for DPO, our ideas can be readily adopted for RLHF based alignment (as these algorithms also require labeled prompt-response pairs). We present the detailed adaption of PROPS for RLHF in Section A.5.
Remark 2 Distinction from Label-DP, Multi-Stage RR and PATE: While the notion of ($\epsilon,\delta$)-Preference Privacy is motivated from the notion of $(\epsilon,\delta)$-Label DP (Chaudhuri & Hsu, 2011; Ghazi et al., 2021a), there are distinctions between the two frameworks. In Label-DP, only labels are treated as private, and noise is added in proportion to the privacy budget to preserve the privacy of the dataset’s labels. In addition, RR (Warner, 1965) is a direct approach to implement preference privacy, as the preference $\ell_{i}$ of a data entry is flipped with probability $\gamma_{\epsilon}=\frac{1}{1+e^{\epsilon}}$. While PROPS shares similarities with Multi-Stage RR Ghazi et al. (2021b) and PATE Papernot et al. (2018), it differs significantly in approach and application. Unlike Multi-Stage RR, which relies on simple sampling for combining noisy labels and model predictions, PROPS uses an MLE-based approach for principled integration. PATE protects privacy for all features using parallel training, whereas PROPS targets preference privacy with a sequential, iterative approach that improves alignment by building on earlier models while preserving privacy.
4 Experiments and Discussion
Datasets and models: In our experiments and validation, we have used (1) three datasets (jondurbin/truthy-dpo-v0.1, Anthropic HH-RLHF, and AlpacaEval) and (2) three different models of varying sizes: Pythia-1B EleutherAI (2024) , GPT2-Large (774M) OpenAI (2024a) and GPT2-Medium (355M) OpenAI (2024b). We have adopted a similar experimental setup as the prior works Rafailov et al. (2024); Chakraborty et al. (2024); von Werra et al. (2020)Lior-Baruch (2024). The code for PROPS is publicly available [^1] https://anonymous.4open.science/r/PROPS-2025.
Evaluations: We provide a comprehensive analysis of the strengths of PROPS in achieving high-quality alignment under privacy constraints. Our results are structured as follows: (a) We compare PROPS with RR across multiple models using the win-tie rate metric, highlighting that PROPS consistently outperforms RR in most cases. (b) We compare PROPS with DP-SGD across various models and datasets to assess its consistent advantage. (c) We also present qualitative examples to show that PROPS can provide better responses than DP-SGD while ensuring the same privacy guarantee. Our results demonstrate the effectiveness of PROPS in delivering privacy-preserving alignment without significant performance degradation.
Table 1: PROPS vs RR based Win-Tie rate on two datasets truthy-dpo-v0.1, AlpacaEval for high-privacy and moderate-privacy regimes with three different models: Pythia-1B, GPT2-Large and GPT2-Medium. In high-privacy regimes, for most of the cases, PROPS outperforms $R$.
AlpacaEval Truthy-DPO
Privacy Budget ($\epsilon$) Pythia GPT2 Large GPT2 Medium Pythia GPT2 Large GPT2 Medium
$0.1$ $52.2$ $46.8$ $55.4$ $66.4$ $61.6$ $72.2$
$0.5$ $64.8$ $75.6$ $86.2$ $56.0$ $71.2$ $60.8$
$1.0$ $59.4$ $70.8$ $84.4$ $63.4$ $52.4$ $46.4$
PROPS vs RR: We summarize the results of comparing PROPS and RR mechanisms in Table 1 using the Win-Tie rate. We implemented a two-stage PROPS ($K=2$), where the first model $M_{1}$ is trained using RR-perturbed labels and the final model $M_{2}$ is trained using PROPS-generated labels. GPT-4 served as the evaluator. We evaluated the performance across three models: Pythia-1B, GPT2-Large, and GPT2-Medium, on the truthy-dpo-v0.1 and AlpacaEval datasets. Results show that PROPS consistently outperforms RR for larger models, especially Pythia-1B. GPT2-Large also outperforms RR in most cases except at $\epsilon=0.1$ on AlpacaEval. GPT2-Medium shows mixed performance, likely due to its limited capacity, leading to occasional underperformance during Stage-2 label generation.
Table 2: PROPS vs DP-SGD based Win-Tie rate on HH-RLHF and truthy-dpo-v0.1 datasets for various privacy budgets, using GPT2-Medium and GPT2-Large models. In high-privacy regimes, PROPS consistently outperforms DP-SGD. Notably, PROPS provides $(\epsilon,0)$-Preference Privacy and DP-SGD provides $(\epsilon,\delta)$-DP where $\delta=10^{-10}$. PROPS simultaneously provides stronger privacy guarantees while leading to better aligned models with higher win-rates (higher utility).
GPT2-Medium GPT2-Large
Privacy Budget ($\epsilon$) $H$-RLHF truthy-dpo $H$-RLHF truthy-dpo
$0.1$ $59.6$ $81.0$ $54.8$ $68.2$
$0.5$ $60.4$ $59.2$ $62.0$ $67.4$
$1.0$ $63.4$ $50.6$ $65.8$ $60.6$
PROPS vs DPSGD: In Table 2, we present Win-Tie rates comparing our proposed algorithm PROPS with the conventional DP-SGD algorithm for GPT2-Large and GPT2-Medium models on the truthy-dpo-v0.1 dataset. DP-SGD was ran using the Gaussian mechanism for 1 epoch with $\delta=10^{-10}$, a gradient clipping threshold of 10, and a batch size of 4 for 2 epochs. As the results indicate, PROPS is able to consistently outperform DP-SGD at higher privacy regimes ($\epsilon=0.1,0.5,1$) for both models. This indicates that while DP-SGD attempts to additionally protect the prompts and responses, it suffers a significant drop in utility for smaller privacy budgets. Additional results are presented in Section A.6. One critical distinction to highlight is that PROPS ensures a pure DP guarantee (i.e. $(\epsilon,0)$-DP) while DP-SGD provides an approximate DP guarantee, denoted as $(\epsilon,\delta)$-DP. We present results for PROPS vs DP-SGD on HH-RLHF, AlpacaEval and truthy-dpo-v0.1 for three privacy parameters. The table indicates that PROPS on-average outperforms DP-SGD at high privacy regimes (Additional results are presented in Section A.6).
Table 3: Win–Tie rate comparison for 2-stage and 3-stage PROPS across privacy budgets on truthy-dpo dataset with GPT2-Large. In high privacy regime ($\epsilon=0.5\ \&\ 1$) 2-stage PROPS outperforms 3-stage PROPS.
PROPS
Privacy Budget ($\epsilon$) 2-stage Wins Ties 3-Stage Wins
$0.5$ $53.2$ $9.2$ $37.6$
$1.0$ $56.8$ $10.4$ $32.8$
$2.0$ $38.4$ $17.6$ $44.0$
Figure 4: Prompt-Response pairs generated by PROPS and DP-SGD based GPT2-Large models and their corresponding scores (helpfulness and harmlessness). The example shows as the privacy constraints become less strict, the quality of responses gradually improves. More prompt-response examples are in Section A.7 of the appendix.
Results on multi-stage PROPS: We now report results to observe if performance gains can be attained by increasing the number of stages used in PROPS. In Table 3, we present results comparing 2-stage and 3-stage PROPS using the truthy-dpo-v0.1 dataset with GPT2-Large. Results indicate that across high privacy budgets, 2-stage PROPS outperforms 3-stage PROPS. This implies that further hyperparameter tuning under privacy constraints is required for 3 or more stages. We leave this as future work as our results up to this point have indicated that the 2-stage results effectively illustrate the key performance trends under varying privacy constraints.
Illustrative Example Responses to Prompts for varying privacy levels: To illustrate the privacy-utility tradeoff, responses from LLMs trained with different privacy levels ($\epsilon=0.1,1,\infty$) were compared as shown in Figure 4. At $\epsilon=0.1$, responses are generic due to high noise; at $\epsilon=1$, models provide more useful but slightly biased answers; and at $\epsilon=\infty$, answers are professional but less helpful. This trend holds across prompts, showing that moderate privacy ($\epsilon=1$) can balance privacy and utility. Response quality was evaluated using GPT-4 based on helpfulness and harmlessness, though final judgments remain subjective. More examples and illustrations of prompt completions under varying privacy budgets are provided in Section A.7.
5 Conclusions
In this paper, we presented new results towards aligning LLMs with Preference-DP, which preserve the privacy of preferences provided by humans. We build and expand upon the concept of label DP for this problem, and present a series of increasingly sophisticated, yet practical privacy preserving mechanisms for alignment. Specifically, starting from a standard randomized response (RR) mechanism which randomly flips human preferences, we presented a new mechanism, PROPS (PROgressively Private Self-alignment) which works across multiple stages. The key insight behind PROPS is that while intermediate LLM models may not yet be fully capable of generating high-quality outputs or responses in the early stages of training, it may still possess sufficient knowledge to correctly label preferences. Thus, our framework leverages the power of intermediate models to enhance alignment efficiency while preserving privacy, offering a novel solution to the challenge of privacy-preserving alignment. We also provided a comprehensive set of experiments on multiple datasets and model sizes which show that PROPS outperforms DP-SGD and randomized response (RR) based approaches. We quantified and measured these gains in terms of Win-Tie rates, and these gains are especially substantial in practically relevant high privacy regimes.
Broader Impact Statement
This work aims to advance the field of privacy-preserving machine learning by addressing the challenge of protecting human labelers’ preferences during the alignment of Large Language Models (LLMs). The proposed PROPS framework ensures alignment quality while maintaining privacy guarantees, which can help mitigate privacy concerns in the growing use of human feedback for training LLMs. By improving the privacy-utility tradeoff, this approach supports the use of such language models while fostering trust in systems that rely on human data. Due to resource constraints, we have used “smaller” language models (e.g. GPT2-Medium, GPT2-Large, Pythia-1B); however, our results still indicate the effectiveness of PROPS in ensuring preference level privacy. Future societal implications of this work include the potential for broader adoption of privacy-preserving methods in language model development, enhancing data security and protecting user identities. However, as there is a need to monitor and ensure that such systems are deployed responsibly and in alignment with ethical guidelines. This work contributes positively to the field by promoting ethical considerations in LLM training and alignment.
Appendix A Appendix
The Appendix is organized as follows:
A.1 Training Details & Comparison of Complexity
A.2 MLE estimator for $\ell^{*}$ using $(\ell_{RR},\ell_{M_{1}})$
A.3 Estimator for $\gamma_{M_{1}}$: Proof of (8)
A.4 Proof Sketch of Theorem 1: Sub-optimality gap for PROPS
A.5 Adapting PROPS for RLHF-based Alignment
A.6 Additional Experimental Results
A.7 Win-rate Evaluation & Additional Prompt-Response Pair Examples as a Function of Privacy Budget
A.1 Training Details & Comparison of Complexity
In this section, we present details on how the models were trained for the experiments. Specifically, our training procedures for each dataset are as follows:
truthy-dpo-v0.1: For this dataset, 15% of the data was used for SFT. The remaining 75% of the data was designated for DPO training. This 75% segment was divided into two halves, with three epochs of DPO run on each half. Subsequently, the dataset was filtered to include only preference pairs generated by prompting a Large Language Model (LLM) to act as ”an honest and helpful assistant.” DPO was then performed on half of this filtered dataset for three epochs. Win-Tie-Loss rates were calculated using the remaining 10% of the Truthy-DPO-v0.1 dataset, which consists of 100 prompts.
HH-RLHF: The HH-RLHF experiment utilitized an existing SFT mode  [^2] https://huggingface.co/jtatman/gpt2-open-instruct-v1-Anthropic-hh-rlhf from Hugging Face that was trained for one epoch on the Anthropic-HH dataset. For DPO, 1000 samples from the test set were used. Specifically, these samples were split into two halves, and DPO was run for three epochs on each half. While the both PROPS and DP-SGD used the same prompts, PROPS receive prompts from the same dataset but in a different format  [^3] https://huggingface.co/datasets/psyche/anthropic-hh-rlhf. Win-Tie-Loss results were generated using 100 samples from the same test set.
AlpacaEval: For AlpacaEval  [^4] https://huggingface.co/datasets/reciprocate/alpaca-eval, 100 examples from the training dataset were used for an initial, quick SFT. Following this, 2,000 examples from the available training dataset were used for DPO training, and 100 examples from the testing dataset were used to evaluate the performance of various DPO methods via Win-Tie-Loss rate. For the PROPS method, the DPO data segment was split in two halves, with three or four epochs of DPO run on each half.
We then provide the training hyperparameters for different models, datasets, and DPO methods in Table 4. For all experiments,
  • DP-SGD based alignment was trained for 1 epoch with a learning rate of 5e55e-5 and batch size of 2.
  • RR based alignment was trained for 3 epochs with a batch size of 4 and a learning rate of $5e^{-5}$, except for Pythia-1B on the Truthy-DPO dataset which was trained with a learning rate of $3e^{-5}$.
  • PROPS based alignment was trained for 2 stages, with each stage using a batch size of 4. A learning rate of $5e^{-5}$ was used in all stages except for GPT2-Large on HH-RLHF and Alpaca, and Pythia-1B on Truthy DPO, where a learning rate of $3e^{-5}$ was used for both stages. Additionally, PROPS was trained for 3 epochs except for the second-stage training of GPT-2 Medium and GPT-2Large on HH-RLHF which was 4 epochs.
Table 4: Overview of hyperparameter configurations for selected DPO methods. Each method is color-coded: green for $R$, teal for PROPS, and yellow for DPSGD.
Method Learning Rate Batch Size Epochs
$R$ $5e^{-5}$ 44 33
PROPS $3e^{-5}$ / $5e^{-5}$ 44 33 / 44
DPSGD $1e^{-5}$ 22 11
Comparison of Complexity In terms of computational complexity (using Big-O notation), the relative costs of each method are as follows:
  • DPSGD: $\mathcal{O}(C_{\text{DPO}})$ — incurs only the standard DPO training cost.
  • $R$: $\mathcal{O}(C_{\text{DPO}}+C_{\text{RR}})$ — includes the DPO cost and cost of applying RR during data processing, which is minimal.
  • PROPS: $\mathcal{O}(C_{\text{DPO}}+C_{\text{RR}}+C_{\text{INF}}+C_{\text{MLE}})$ — combines the DPO cost, RR overhead, model inference before the second stage, and the cost of maximum likelihood estimation (MLE), which is minimal.
DPSGD is the most computationally efficient, involving only the cost of standard DPO optimization. RR adds moderate overhead due to randomly flipping preferences, but is still lightweight. PROPS incurs the highest computational cost, as it integrates several components: the baseline DPO cost, the randomized response mechanism, model inference over a second data partition, and an additional phase of maximum likelihood estimation (MLE).
A.2 MLE estimator for $\ell^{*}$ using $(\ell_{RR},\ell_{M_{1}})$
Given the flipped labels $\ell_{RR}$ and $\ell_{M_{1}}$ by RR and model $M_{1}$, respectively, we aim to come up with a good decision-making policy for the proposed algorithm. We calculate the likelihood of observing $(\ell_{RR},\ell_{M_{1}})$ given the possible values of $\ell^{*}$. We define $\gamma_{\epsilon}$ as the flipping probability of RR and $\gamma_{M_{1}}$ as the flipping probability of the model.
Figure 5: The table represents the probability of observing $\ell_{RR}$ and $\ell_{M}$ based on the flipping probabilities $\gamma_{\epsilon}$ and $\gamma_{M_{1}}$ and true label $\ell^{*}$.
For binary $\ell^{*}$, now we present the probability of observing specific values of $\ell_{RR}$ and $\ell_{M_{1}}$. To find the best estimator, we compute the log-likelihood ratio:
$\Lambda$ $=\log\left(\frac{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=0)}{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=1)}\right)$
$\overset{(a)}{=}\log\left(\frac{\mathbb{P}(\ell_{RR}|\ell^{*}=0)\cdot\mathbb{P}(\ell_{M_{1}}|\ell^{*}=0)}{\mathbb{P}(\ell_{RR}|\ell^{*}=1)\cdot\mathbb{P}(\ell_{M_{1}}|\ell^{*}=1)}\right)$
$\overset{(b)}{=}(-1)^{\ell_{RR}}\log\left(\frac{1-\gamma_{\epsilon}}{\gamma_{\epsilon}}\right)+(-1)^{\ell_{M_{1}}}\log\left(\frac{1-\gamma_{M_{1}}}{\gamma_{M_{1}}}\right)$
where, (a) is obtained since $\ell_{RR}$ and $\ell_{M_{1}}$ are independent, and (b) follows from using the expressions derived in Table 5.
A.3 Estimator for $\gamma_{M_{1}}$: Proof of (8)
We have noisy labels $\ell_{RR}$ generated by RR with a flipping probability of $\gamma_{\epsilon}$, and predicted labels $\ell_{M_{1}}$ by the model $M_{1}$ with a flipping (error) probability of $\gamma_{M_{1}}$. We define $\ell_{RR}$ and $\ell_{M_{1}}$ as:
$\ell_{RR}=\ell^{*}\oplus U,\quad\ell_{M_{1}}=\ell^{*}\oplus V,$ (9)
where $U\sim\text{Bernoulli}(\gamma_{\epsilon})$ and $V\sim\text{Bernoulli}(\gamma_{M_{1}})$. We first make the observation that for $i^{th}$ sample in the dataset, $\ell_{M_{1}}^{(i)}\oplus\ell_{RR}^{(i)}=(\ell^{*}_{i}\oplus V_{i})\oplus(\ell^{*}_{i}\oplus V_{i})=V_{i}\oplus U_{i}$, where, $U_{i}$ and $V_{i}$ are independent. Now, define
$\mu_{M_{1}}$ $=\frac{\sum_{i}\ell^{(i)}_{M_{1}}\oplus\ell^{(i)}_{RR}}{|\mathcal{D}_{2}|}$
and note that $\mu_{M_{1}}$ is an unbiased estimator for the expected value of $\mathbb{E}[U\oplus V]=\gamma_{M_{1}}(1-\gamma_{\epsilon})+\gamma_{\epsilon}(1-\gamma_{M_{1}}))$. Hence, we can use $\mu_{M}$ to obtain an estimate for $\hat{\gamma}_{M_{1}}$ as follows:
$\hat{\gamma}_{M_{1}}$ $=\frac{(\mu_{M_{1}}-\gamma_{\epsilon})}{1-2\gamma_{\epsilon}}$ (10)
This concludes the proof of equation 8. To mitigate potential correlation when estimating $\gamma_{M_{1}}$, we employ disjoint datasets across the two stages of PROPS. In stage-1, $M_{1}$ is aligned using one half of the dataset, ensuring that the learned parameters are independent of the data used in subsequent steps. In stage-2, MLE with RR and model-generated labels from $M_{1}$ on the remaining half of the dataset are used to generate a more aligned model $M_{2}$. The use of disjoint data across stages reduces the risk of direct dependence between the two datasets.
Validation of our estimation procedure for $\gamma^{*}_{M_{1}}$: To validate the robustness of our estimation process for obtaining $\hat{\gamma_{M_{1}}}$, we compare the estimated error rate of model $M_{1}$ with its true counterpart (the ”oracle” error rate of $\gamma_{M_{1}}^{*}$). Specifically, Table 5 presents a comparison between the estimated error rate $\hat{\gamma}_{M_{1}}$, computed using equation 8 from the paper, and the oracle error rate $\hat{\gamma}_{M_{1}}^{*}$, obtained by evaluating $M_{1}$’s predictions on the unperturbed preference data against the ground truth (i.e., the original preferences prior to flipping). The results demonstrate that, across the high privacy regime, the estimated and oracle error rates are consistently well-aligned, indicating that our estimation process is accurate and reliable even under strict privacy constraints.
Table 5: Comparable estimations on flipping for our method ($\hat{\gamma_{M_{1}}}$) compared to the model flipping probability obtained from the ”oracle” setting $\gamma_{M_{1}}^{*}$.
Privacy Budget $(\epsilon)$ $\hat{\gamma_{M_{1}}}$ (Our Method) $\gamma^{*}_{M_{1}}$ (Obtained from Oracle)
5 0.277 0.268
2 0.334 0.362
1 0.415 0.366
0.5 0.468 0.402
0.1 0.421 0.413
A.4 Proof Sketch of Theorem 1: Sub-optimality gap for PROPS
Chowdhury et al. (2024a) provides a bound on the sub-optimality gap between the rewards obtained using an optimal model aligned under DPO with noiseless preference data and a model aligned under their proposed robust DPO (rDPO) method, which accounts for noisy (i.e. flipped) preference data. Their result characterizes how many preference samples are needed at different noise levels to ensure that the loss in rewards (relative to the original DPO method) does not exceed a certain bound. Chowdhury et al. (2024a) assume that the model can be expressed as a log-linear policy (i.e. as a function of a feature map $\phi(x,y)^{T}$ and the weights of the last layer $\theta$). They assume that the policy (characterized by the parameters $\theta$) and its first and second order gradients are bounded, to provide bounds and Lipschitz guarantees on the difference between rewards attained for the chosen and rejected responses. It is also assumed that the parameters $\theta$ are in the following set: $\{\theta\in\mathbb{R}^{d}|\sum_{i=1}^{d}\theta_{i}=0\}$. Additionally, they assume that the fine-tuned model (i.e. model before alignment) has a good coverage of the feature space, to ensure that the relative condition number $\kappa$ between the covariance matrices of the aligned and fine-tuned policies is small. They prove that their sub-optimality gap is as follows:
$\mathcal{O}\left(\frac{\sqrt{\kappa}}{\gamma\beta(1-2\epsilon)}\sqrt{\frac{d}{n}}\right),$ (11)
where $\epsilon$ denotes the flipping rate, $n$ represents the number of samples used for training, $d$ denotes the dimension of the features space, and $\gamma$ is a constant that depends on $\beta$ and the bound on $\theta$. We leverage the above assumptions to derive a bound on the sub-optimality gap between the optimal non-private DPO policy parameters and the policy obtained through two-stage PROPS.
Deriving a sub-optimality gap for PROPS requires knowing how often a preference label is flipped during stage 2. Recall that during stage 2 of PROPS, a partially aligned model $M_{1}$ predicts the preference labels of prompt-response pairs from $D_{2}$, denoted as $l_{M_{1}}$. The flipping rate of $M_{1}$ is represented by its error rate $\gamma_{M_{1}}$. These labels are then combined with labels from $D_{2}$ obtained with RR, where the labels $l_{RR}$ are flipped with rate $\gamma_{\epsilon}=\frac{1}{1+e^{\epsilon}}$, via an MLE to obtain labels $l_{\text{PROPS}}$. These final labels are then used to align $M_{1}$ to obtain model $M_{2}$.
To analyze the label flipping probability of MLE ($\gamma_{\text{MLE}}$), we derive all possible scenarios where the MLE flips the preference labels, and we assume $\gamma_{M_{1}}<\gamma_{\epsilon}\leq\frac{1}{2}$. Recall that in Section A.2, the log-likelihood of observing labels flipped by RR $\ell_{RR}$ and labels generated by the partially aligned model $M_{1}$ (denoted as $\Lambda(\ell_{RR},\ell_{M_{1}}$)) for binary ground truth preferences $\ell^{*}$ can be expressed as follows:
$\small\Lambda(\ell_{RR},\ell_{M_{1}})$ $=\log\left(\frac{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=0)}{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=1)}\right)$
$=(-1)^{\ell_{RR}}\cdot\log\left(\frac{1-\gamma_{\epsilon}}{\gamma_{\epsilon}}\right)+(-1)^{\ell_{M_{1}}}\cdot\log\left(\frac{1-\gamma_{M_{1}}}{\gamma_{M_{1}}}\right).$
Therefore, PROPS can generate label $\ell_{\text{PROPS}}$ for each prompt as:
$\!\ell_{{\text{PROPS}}}\!(\ell_{RR},\!\ell_{M_{1}})\!=\!\begin{cases}1,\quad\text{if }\Lambda(\ell_{RR},\!\ell_{M_{1}})\!\leq\!0\\ 0,\quad\text{if }\Lambda(\ell_{RR},\!\ell_{M_{1}})\!>\!0.\end{cases}$
We now provide an overview of the decisions that the Maximum Likelihood Estimator (MLE) can make based on the possible combinations of $(\ell_{RR},\ell_{M_{1}})$.
Case 1: ($\ell_{RR}=\ell_{M_{1}}=0$)
In this case we can compute $\Lambda(\ell_{RR},\ell_{M_{1}})=\log\left(\frac{(1-\gamma_{\epsilon})\cdot(1-\gamma_{M_{1}})}{\gamma_{\epsilon}\cdot\gamma_{M_{1}}}\right)>0$, therefore $\ell_{\text{PROPS}}(\ell_{RR},\ell_{M_{1}})=0$.
Case 2: ($\ell_{RR}=\ell_{M_{1}}=1$)
In this case we can compute $\Lambda(\ell_{RR},\ell_{M_{1}})=\log\left(\frac{\gamma_{\epsilon}\cdot\gamma_{M_{1}}}{(1-\gamma_{\epsilon})\cdot(1-\gamma_{M_{1}})}\right)\leq 0$, therefore $\ell_{\text{PROPS}}(\ell_{RR},\ell_{M_{1}})=1$.
Case 3: ($\ell_{RR}=1,\ell_{M_{1}}=0$)
For this scenario $\Lambda(\ell_{RR},\ell_{M_{1}})=\log\left(\frac{\gamma_{\epsilon}\cdot(1-\gamma_{M_{1}})}{(1-\gamma_{\epsilon})\cdot\gamma_{M_{1}}}\right)>0$, therefore $\ell_{\text{PROPS}}(\ell_{RR},\ell_{M_{1}})=0$.
Case 4: ($\ell_{RR}=0,\ell_{M_{1}}=1$)
For this scenario $\Lambda(\ell_{RR},\ell_{M_{1}})=\log\left(\frac{\gamma_{M_{1}}\cdot(1-\gamma_{\epsilon})}{(1-\gamma_{M_{1}})\cdot\gamma_{\epsilon}}\right)\leq 0$, therefore $\ell_{\text{PROPS}}(\ell_{RR},\ell_{M_{1}})=1$.
In each of the four possible scenarios, we observe that for $\gamma_{M_{1}}<\gamma_{\epsilon}$, the predictions of the MLE match the predictions of $M_{1}$. This implies that $\gamma_{\text{MLE}}=\gamma_{M_{1}}$. Conversely, when $\gamma_{\epsilon}<\gamma_{M_{1}}$, performing the same analysis shows that the MLE will match the predictions of RR, resulting in $\gamma_{\text{MLE}}=\gamma_{\epsilon}$. Therefore, the flipping rate of PROPS can be described as $\gamma_{\text{MLE}}=\min(\gamma_{\epsilon},\gamma_{M_{1}})$. Thus, using the smoothness assumptions and bound provided by Chowdhury et al. (2024a), the sub-optimality gap between between the optimal non-private DPO policy parameters $\theta^{*}$ and the policy parameters obtained through two-stage PROPS $\hat{\theta}_{\text{PROPS}}$ can be obtained as:
$\small\underbrace{\left\lVert\hat{\theta}_{\text{PROPS}}-\theta^{*}\right\rVert}_{\text{Sub-Optimality Gap}}\leq\mathcal{O}\left(\frac{\sqrt{\kappa}}{\gamma\beta(1-2\cdot\min(\gamma_{M_{1}},\gamma_{\epsilon}))}\sqrt{\frac{d}{n_{2}}}\right),$
where $n_{2}$ is the number of samples used in stage 2 of PROPS.
A.5 Adapting PROPS for RLHF-based Alignment
PROPS can be effectively adapted for Reinforcement Learning with Human Feedback (RLHF)-based alignment, enhancing privacy preservation without compromising performance. In RLHF, a reward model is first trained on a preference dataset, which is then used to optimize a fine-tuned model via Proximal Policy Optimization (PPO). To ensure preference privacy in RLHF, PROPS can be adapted in a multi-stage framework as follows, demonstrated here for a 2-stage setup:
  1. 1.
    Dataset Partitioning: Divide the preference dataset into two disjoint subsets, $D_{1}$ and $D_{2}$, ensuring data privacy and enabling staged alignment.
  2. 2.
    Training on $D_{1}$: Apply randomized response (RR) on $D_{1}$ to protect preference privacy. Use the perturbed data to train a reward model and partially align the fine-tuned model via PPO, resulting in a preliminary model, $M_{1}$.
  3. 3.
    Ranking with $M_{1}$: Apply RR on $D_{2}$ to privatize preferences in this subset. Then, use $M_{1}$ to rank the responses in $D_{2}$.
  4. 4.
    Combining Rankings for Label Generation: Combine the rankings derived from RR-perturbed preferences on $D_{2}$ with the rankings provided by $M_{1}$. Use a maximum-likelihood estimation (MLE) approach to generate new, privacy-preserving labels for $D_{2}$ based on these combined rankings.
  5. 5.
    Updating the Reward Model: Use the new labels from $D_{2}$ to update the reward model, which then produces a more refined, partially aligned model, $M_{2}$.
This staged approach ensures that each model leverages prior knowledge while progressively refining alignment in a privacy-preserving manner. By iteratively updating the reward model and partially aligning the fine-tuned model, PROPS achieves an optimal balance between privacy and alignment quality. The method can be extended to more stages as needed, providing flexibility and scalability for RLHF-based alignment tasks.
Table 6: PROPS vs RR based Win-Tie rate on two datasets truthy-dpo-v0.1, AlpacaEval for high-privacy and moderate-privacy regimes with three different models: Pythia-1B, GPT2-Large and GPT2-Medium. In high-privacy regimes, PROPS consistently outperforms $R$.
AlpacaEval Truthy-DPO
Privacy Budget ($\epsilon$) Pythia GPT2 Large GPT2 Medium Pythia GPT2 Large GPT2 Medium
$0.1$ $52.2\pm 4.26$ $46.8\pm 3.31$ $55.4\pm 1.62$ $66.4\pm 3.44$ $61.6\pm 1.01$ $72.2\pm 4.62$
$0.5$ $64.8\pm 6.79$ $75.6\pm 3.87$ $86.2\pm 2.4$ $56.0\pm 4.24$ $71.2\pm 2.85$ $60.8\pm 5.26$
$1.0$ $59.4\pm 3.87$ $70.8\pm 3.42$ $84.4\pm 2.8$ $63.4\pm 4.96$ $52.4\pm 4.71$ $46.4\pm 4.22$
$2.0$ $51.0\pm 3.40$ $37.6\pm 3.07$ $75.4\pm 3.77$ $62.2\pm 5.81$ $58.0\pm 3.03$ $54.8\pm 0.74$
Table 7: PROPS vs DP-SGD based Win-Tie rate on HH-RLHF, truthy-dpo-v0.1 datasets for high-privacy and moderate-privacy regimes on GPT2-Medium, and GPT2-Large models. As we can observe, in high-privacy regime, PROPS consistently outperforms DP-SGD.
GPT2-Medium GPT2-Large
Privacy Budget ($\epsilon$) $H$-RLHF truthy-dpo $H$-RLHF truthy-dpo
$0.1$ $59.6\pm 6.08$ $81.0\pm 8.41$ $54.8\pm 4.62$ $68.2\pm 5.52$
$0.5$ $60.4\pm 3.00$ $59.2\pm 4.44$ $62.0\pm 3.22$ $67.4\pm 6.08$
$1.0$ $63.4\pm 7.22$ $50.6\pm 4.63$ $65.8\pm 5.19$ $60.6\pm 4.31$
$2.0$ $45.4\pm 5.08$ $61.2\pm 7.54$ $63.8\pm 4.87$ $46.6\pm 5.12$
Table 8: Win–Tie rate comparison for 2-stage and 3-stage PROPS across privacy budgets on truthy-dpo dataset with GPT2-Large. In high privacy regime ($\epsilon=0.5\ \&1$) 2-stage PROPS outperforms 3-stage PROPS.
PROPS
Privacy Budget ($\epsilon$) 2-stage Wins Ties 3-Stage Wins
$0.5$ $53.2\pm 3.35$ $9.2\pm 2.28$ $37.6\pm 3.58$
$1.0$ $56.8\pm 9.12$ $10.4\pm 1.67$ $32.8\pm 8.67$
$2.0$ $38.4\pm 7.27$ $17.6\pm 3.58$ $44.0\pm 10.58$
A.6 Additional Experimental Results
In this section, we provide more context and analysis on the results presented in the main paper.
PROPS vs DP-SGD and PROPS vs RR: We first present supplementary experimental results and their corresponding standard derivations. Specifically, we report means and standard deviations of the win-tie rates of PROPS vs. RR and PROPS vs. DP-SGD in Tables 6 and 7 respectively. Unlike DP-SGD, which employs gradient perturbation, PROPS utilizes an input perturbation mechanism that maintains the post-processing property of differential privacy. This inherent flexibility enables extensive hyper-parameter tuning, including training epochs, without compromising privacy guarantees. Consequently, despite fluctuations in the win-tie-loss ratio as determined by the GPT-4 model, we still observe consistent trends. Notably, in the high privacy regime (i.e., a low privacy budget), PROPS exhibits a higher win rate. Conversely, in scenarios with a more relaxed privacy constraint, DP-SGD demonstrates better performance. Another noteworthy discrepancy arises from the interplay between model size and model family as larger models benefit from PROPS more. This discrepancy may be attributed to the different learning capacities of the model under varying privacy requirements. Specifically, when the preference dataset is subjected to consistent noise injection, the larger models in the initial training stages exhibit better learning, thereby positively influencing subsequent model iterations.
However, the advantage conferred by larger models diminishes when the initial models fail to adequately align with human values as the preference dataset is subject to huge noise injection. Thus, GPT-2 Large models exhibit superior performance compared to GPT-2 Medium models, owing to their enhanced learning capabilities within the PROPS framework. More evidently, as shown in the win-tie-loss rate compassion between different datasets, when the preference dataset becomes more complex, models in the initial stages of PROPS fail to guide and provide extra help for subsequent stages.
Additional Results on Multi-stage PROPS. We present the mean and standard deviation of 2-Stage vs 3-Stage PROPS with GPT2-Large model on truthy-dpo dataset for 5 consecutive trials in Table 8. This suggests that further hyperparameter tuning is necessary for three or more stages when operating under privacy constraints. The large standard deviation indicates that performance can vary significantly, likely due to the impact of these parameters. Nevertheless, we conclude that for high-privacy regimes ($\epsilon=0.5\ \&1$), a 2-stage approach is preferable, and a thorough study of hyperparameter selection is still required.
A.7 Win-rate Evaluation & Additional Prompt-Response Pair Examples as a Function of Privacy Budget
In this section we explain how win-tie rate measurements were obtained for the experiments and show additional prompt-response pair examples at different privacy budgets.
Win-rate Evaluation: We used GPT4 to obtain win-tie rate measurements by comparing responses generated by PROPS against conventional privacy-based alignment methods (DP-SGD and RR). Figure 6 shows an example of how these measurements were obtained. We instruct GPT4 to act as an evaluator/judge (to assign the Helpfulness and Harmlessness scores) by providing it the following prompt similar to Chakraborty et al. (2024), “You are a helpful, harmless, and precise assistant for checking the quality of the answer. We would like to request your feedback on the performance of two AI assistants in response to the user question. Please rate the harmlessness, helpfulness, and level of detail of their responses. Your evaluation should consider factors such as the helpfulness, harmlessness, relevance, accuracy, depth, creativity, and level of detail of the response. Bear in mind that the response might be truncated at the end due to length constraints. Each assistant receives an overall score on a scale of 1 to 10 (or 1 to 5), where a higher score indicates better overall performance. Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.” For the same prompt, “Do violent video games cause people to become more violent in real life”, as shown in Figure 6, we provide GPT4 responses generated by GPT2-Large trained under PROPS and DP-SGD. As the figure shows, GPT4 assigns a higher score to the PROPS response, ultimately giving it a win for this prompt.
Figure 6: Illustration of using GPT4 to evaluate overall quality of responses generated by a model trained via PROPS and DP-SGD.
Illustrative Example Responses to Prompts for varying privacy levels for PROPS and DP-SGD based alignment: In this section, we present a detailed evaluation of prompt-response pairs generated by a GPT2-Large, and GPT2-Medium models aligned using two different privacy-preserving mechanisms: PROPS and DP-SGD. These pairs were created under varying privacy budgets to investigate how alignment strategies perform under different levels of privacy constraints. Figure 7 offers a comparative analysis of the responses generated by the two approaches across identical prompts. Each prompt was used to elicit model outputs under multiple privacy regimes, allowing us to evaluate the trade-offs between privacy, harmlessness, and helpfulness in generated content.
Figure 7: Prompt-Response pairs generated by GPT2-Large model based on PROPS and DP-SGD alignment for different privacy regimes.
Our analysis reveals a consistent pattern across both mechanisms: a strong preference for harmlessness over helpfulness. That is, the responses tend to prioritize minimizing potential harm, often at the cost of informativeness. This behavior aligns with a fundamental goal of language model alignment—to ensure that model outputs remain consistent with human ethical and safety standards. However, there is a notable difference in how each method balances these trade-offs. For PROPS, we observe a positive correlation between increasing privacy budgets and the helpfulness of responses, suggesting that it can maintain utility while still adhering to privacy constraints. In contrast, DP-SGD does not exhibit a clear trend in helpfulness improvement with higher privacy budgets, indicating potential limitations in its ability to retain utility under stronger privacy guarantees. These findings highlight that PROPS may offer a more favorable approach when seeking to balance privacy preservation with the utility of model outputs in alignment tasks.
Figure 9 shows two prompt-response examples generated by PROPS, DP-SGD, and RR. Similar trends are observed between PROPS and DP-SGD as in the previous examples, where DP-SGD generates similar responses across privacy budgets while PROPS improves in response quality when the privacy budget is increased. RR also seems to exhibit similar behaviors where for the prompt, “Can humans catch warts from frogs or toads?”, it generates a more focused response at $\epsilon=0.1$ rather than $\epsilon=1$. However, looking at $\epsilon=0.1$, the figure indicates that PROPS is able to generate an answer that better addresses the concern of the user compared to vanilla RR.
Figure 8: Prompt-Response pairs generated by GPT2-Large and GPT2-medium models based on PROPS and DP-SGD alignment for different privacy regimes.
Figure 9: Prompt-Response pairs generated by GPT2-Large based on PROPS, DP-SGD, AND $R$ alignment for different privacy regimes.

### v2

@@
 The process of aligning LLMs relies on datasets comprising prompts, LLM-generated responses, and preference labels that indicate which response aligns better with human values, collectively referred to as preference data.
-The alignment approaches, including reinforcement learning with human feedback (RLHF) (Stiennon et al., 2020) and direct preference optimization (DPO) (Rafailov et al., 2024), leverage these preference datasets with ranked labels provided by human annotators.
+The alignment approaches, including reinforcement learning with human feedback (RLHF) (stiennon2020learning) and direct preference optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage), leverage these preference datasets with ranked labels provided by human annotators.
 Figure 1: (a) Randomized Response (RR) based alignment where human preferences in the dataset $\mathcal{D}$ are privatized using $R$ which are then used for alignment.
@@
 These progressively refined private preferences are then used for alignment to arrive at the final model $M_{2}$.
-Motivation for Human Preference Privacy: Relying on human feedback for alignment brings up privacy concerns, particularly around the extent to which a labeler’s preferences might unintentionally expose their underlying values, identity, or personality.
-In medical applications, for example, publicly available case reports or anonymized symptoms can be used to prompt LLMs to generate diagnoses, with human feedback aligning these outputs to clinical best practices (Reddy, 2023; Han et al., 2024).
-Here, the preference data represents expert judgment about nuanced medical decisions and must remain private to protect the integrity and confidentiality of clinical expertise.
-In Wang et al.
-(2024), a large number of consultation records were used to construct a dataset to fine-tune an LLM to act as a chatbot physician, with GPT-4 used to retain records that highlighted professionalism, explainability, and emotional support.
+Motivation for Human Preference Privacy: While preference data can significantly improve the alignment of large models with expert reasoning, it often carries deep privacy risks—particularly in domains where human feedback encodes sensitive strategies, values, or professional heuristics.
+This tension is most evident in several high-stakes application scenarios as we discuss next.
+In clinical decision support systems, for example, physicians’ feedback reflects diagnostic reasoning and treatment heuristics tied to protected health information and institutional IP; disclosure can undermine both patient privacy and clinical competitiveness (reddy2023evaluating; han2024towards).
+In legal and judicial settings, alignment with lawyers’ or judges’ feedback risks revealing privileged deliberations, litigation strategies, or interpretive biases that must remain confidential to preserve due process.
+These scenarios illustrate a central point: preference privacy is not merely a theoretical consideration but a critical requirement for deploying aligned models in the most sensitive and societally consequential domains.
+The privacy–utility trade-offs in such settings demand specialized alignment mechanisms that protect individual and institutional preference signals while preserving their utility for model improvement.
+In wang2024gpt, a large number of consultation records were used to construct a dataset to fine-tune an LLM to act as a chatbot physician, with GPT-4 used to retain records that highlighted professionalism, explainability, and emotional support.
 As the model was shown to be effective, this indicates that certain preferences of a doctor’s decision-making can be identified which could lead to potential exposure of a doctor’s preferences.
-Also, in policy analysis (Rao et al., 2023), publicly available survey questions or proposals can be used to elicit LLM-generated analyses, where policymakers’ feedback reveals sensitive interpretative insights that may require protection.
-Bakker et al.
-(2021), for example, observed that placing more emphasis on politics in their surveys to participants “resulted in self-reports of personality traits that were in some cases more aligned with preexisting political preferences.”
+Also, in policy analysis (rao2023ethical), publicly available survey questions or proposals can be used to elicit LLM-generated analyses, where policymakers’ feedback reveals sensitive interpretative insights that may require protection.
+bakker2021reconsidering, for example, observed that placing more emphasis on politics in their surveys to participants “resulted in self-reports of personality traits that were in some cases more aligned with preexisting political preferences."
 Recent work on privacy-preserving fine-tuning and alignment (see Section 2) typically treats prompts, responses, and human feedback as jointly private.
@@
 Preference level privacy ensures that for any tuple $(x,y_{1},y_{2})$, where $x$ denotes the prompt and $y_{1},y_{2}$ denote the LLM-generated responses, the individual human-preference $\ell^{*}$ (which denotes whether $y_{1}$ or $y_{2}$ is preferable) does not significantly impact the aligned model.
-Formally, we leverage the existing notion of Differential Privacy (DP) Dwork et al.
-(2014), and use it to formalize the notion of $(\epsilon,\delta)$-preference-level DP, where $(\epsilon,\delta)$ represent the privacy budgets.
+Formally, we leverage the existing notion of Differential Privacy (DP) dwork2014algorithmic, and use it to formalize the notion of $(\epsilon,\delta)$-preference-level DP, where $(\epsilon,\delta)$ represent the privacy budgets.
 The notion of labeler-level privacy (also commonly referred to as user-level privacy in the DP literature) hides the presence/absence of any individual human labeler and protects all the labels annotated by the labeler.
@@
 This staged process leverages intermediate models to improve preference labeling and reduce reliance on noisy labels, enhancing alignment quality while maintaining privacy.
-As the model evolves over $K$ stages, alignment becomes progressively refined.
 PROPS effectively balances privacy preservation with performance, offering a novel alignment framework.
   • Theoretical Insights.
-We study the utility-privacy tradeoffs for PROPS by analyzing the Sub-optimality gap Chowdhury et al.
-(2024a) defined as the difference between between the weights of a non-privately trained model with the weights of a privately trained model with a privacy budget $\epsilon$ in Section 3.1 (Theorem 1).
+We study the utility-privacy tradeoffs for PROPS by analyzing the Sub-optimality gap chowdhury2024provably defined as the difference between between the weights of a non-privately trained model with the weights of a privately trained model with a privacy budget $\epsilon$ in Section 3.1 (Theorem 1).
 The upper bound on this gap shows that PROPS (which uses MLE combining in the second stage) is no worse than vanilla RR and performs better as long as the intermediate model gets better at predicting preferences.
@@
 (b) Prompt-Response pairs generated by GPT2-Large model with PROPS, DP-SGD and $R$-based alignment for different privacy regimes.
+Consistent with standard approaches for alignment, this work focuses on the common setting of binary preferences (pairwise comparisons).
+The core ideas could potentially be extended to multiple preferences using techniques in Zhu2023Principled, we leave this generalization as future work.
 Related works & Limitations: We next provide an overview of LLM alignment and the associated privacy risks in using human preference data.
@@
 The alignment process relies on a dataset $\mathcal{D}$ consisting of $n$ samples, each containing a prompt $x$, LLM-generated responses $(y_{1},y_{2})$, and a human-preferred label $\ell^{*}$, collectively referred to as preference data.
-Two conventional methods for alignment, Reinforcement Learning with Human Feedback (RLHF) (Stiennon et al., 2020) and Direct Preference Optimization (DPO) (Rafailov et al., 2024), utilize preference datasets with ranked labels provided by human annotators.
+Two conventional methods for alignment, Reinforcement Learning with Human Feedback (RLHF) (stiennon2020learning) and Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage), utilize preference datasets with ranked labels provided by human annotators.
 While these aligned models improve the quality of generated responses, they introduce privacy risks, particularly concerning the identity of the annotators and their associated preferences.
-Recent work shows membership inference attack on preference data for LLM alignment (Feng et al., 2024), and highlights the vulnerabilities of using human annotated preference data during alignment.
-To mitigate the privacy risks associated with human-annotated preference data, the notion of differential privacy (DP) (Dwork et al., 2014) has recently been explored for fine-tuning and alignment of LLMs.
-For example, Yu et al.
-(2021) applied DP to fine-tuning by introducing privacy guarantees for smaller, appended parameters such as LoRA and adapters.
-Singh et al.
-(2024) introduced a two-stage fine-tuning process, and Yu et al.
-(2024) addressed the privacy-preserving alignment challenge by ensuring DP protection for users’ prompts against labelers during the generation of preference datasets for alignment.
-Additionally, Wu et al.
-(2023) proposed applying DP to RLHF by splitting the dataset into three disjoint sets to ensure DP at each stage of RLHF.
-Moreover, Feng et al.
-(2024) investigated the vulnerability of LLMs aligned using human preference datasets to membership inference attacks (MIAs), and provided empirical evidence that DPO models are more vulnerable to MIA compared to RLHF based models.
-Private Alignment: To address privacy threats, the idea of Differential Privacy (DP) (Dwork et al., 2014) has been integrated into the LLM training process, including pre-training, fine-tuning, and alignment, to protect the privacy of preference data.
-To achieve DP, current state-of-the-art approaches utilize DP-SGD (Abadi et al., 2016), which privatizes gradients to ensure the privacy of the entire preference data throughout the training process.
-Previous literature has explored key aspects relevant to this study: (Chowdhury et al., 2024b) examined privacy preserving reward estimation for RLHF methods, while (Chowdhury et al., 2024a) investigated the robustness of DPO with noisy preference datasets, providing the foundational basis for this study.
-However, to the best of our knowledge, the concept of Label Differential Privacy (Label-DP) (Ghazi et al., 2021b) has not been explored in the context of alignment.
+Recent work shows membership inference attack on preference data for LLM alignment (feng2024exposing), and highlights the vulnerabilities of using human annotated preference data during alignment.
+To mitigate the privacy risks associated with human-annotated preference data, the notion of differential privacy (DP) (dwork2014algorithmic) has recently been explored for fine-tuning and alignment of LLMs.
+For example, yu2021differentially applied DP to fine-tuning by introducing privacy guarantees for smaller, appended parameters such as LoRA and adapters.
+singh2024whispered introduced a two-stage fine-tuning process, and yu2024privacy addressed the privacy-preserving alignment challenge by ensuring DP protection for users’ prompts against labelers during the generation of preference datasets for alignment.
+Additionally, wu2023privately proposed applying DP to RLHF by splitting the dataset into three disjoint sets to ensure DP at each stage of RLHF.
+Moreover, feng2024exposing investigated the vulnerability of LLMs aligned using human preference datasets to membership inference attacks (MIAs), and provided empirical evidence that DPO models are more vulnerable to MIA compared to RLHF based models.
+Private Alignment: To address privacy threats, the idea of Differential Privacy (DP) (dwork2014algorithmic) has been integrated into the LLM training process, including pre-training, fine-tuning, and alignment, to protect the privacy of preference data.
+To achieve DP, current state-of-the-art approaches utilize DP-SGD (abadi2016deep), which privatizes gradients to ensure the privacy of the entire preference data throughout the training process.
+Previous literature has explored key aspects relevant to this study: (chowdhury2024differentially) examined privacy preserving reward estimation for RLHF methods, while (chowdhury2024provably) investigated the robustness of DPO with noisy preference datasets, providing the foundational basis for this study.
+However, to the best of our knowledge, the concept of Label Differential Privacy (Label-DP) (ghazi2021deep) has not been explored in the context of alignment.
 Label-DP could enable models to achieve high utility while reducing privacy leakage, particularly when protecting the labels generated by human annotators.
@@
 We note that the problem of preference-level privacy is similar to the problem of robust alignment in the presence of noisy preferences.
-Specifically, recent works, including Mitchell (2023), Chowdhury et al.
-(2024a) and Chowdhury et al.
-(2024b) study the robustness of alignment when the human-annotated labels are intrinsically noisy.
+Specifically, recent works, including cdpo, chowdhury2024provably and chowdhury2024differentially study the robustness of alignment when the human-annotated labels are intrinsically noisy.
 The distinction, however, is the following: in our setting, the injected noise (and more importantly, the parameters of the privacy-preserving mechanisms (detailed in the next Section)) are known and can be controlled as a function of the privacy parameters.
+Our framework can be broadly positioned within the literature on self-training, where a model ($M_{1}$) generates labels for a subsequent model ($M_{2}$).
+However, our core novelty is distinct.
+Standard self-training does not operate in a privacy-preserving context.
+Our contribution lies in the mechanism for combining a known privacy mechanism (RR) with the unknown-quality predictions of the $M_{1}$ model.
+The novelty is our method for privately estimating the intermediate model’s error rate ($\hat{\gamma}_{M_{1}}$) using the disagreement between the two noisy signals, which allows us to provably create a higher-quality private label set than using either signal alone.
 2 Preliminaries on Alignment & Privacy
@@
 The instant-specific true loss mentioned in Equation equation 1 represents the loss for every prompt $x$ in preference data $\mathcal{D}$, therefore the expected DPO loss can be defined as:
-$\mathbb{E}[\textsf{loss}(x,y_{1},y_{2},\ell^{*})]$ $=\mathbb{E}_{(x,y_{1},y_{2},\ell^{*})\sim\mathcal{D}}\big{\{}\mathds{1}(\ell^{*}=1)\cdot\textsf{loss}(x,y_{1}\succ y_{2})$
-$+\mathds{1}(\ell^{*}=0)\cdot\textsf{loss}(x,y_{2}\succ y_{1})\big{\}}.$ (2)
+$\mathbb{E}[\textsf{loss}(x,y_{1},y_{2},\ell^{*})]=\mathbb{E}_{(x,y_{1},y_{2},\ell^{*})\sim\mathcal{D}}\big\{\mathds{1}(\ell^{*}=1)\cdot\textsf{loss}(x,y_{1}\succ y_{2})+\mathds{1}(\ell^{*}=0)\cdot\textsf{loss}(x,y_{2}\succ y_{1})\big\}.$ (2)
 2.1 Privacy for Alignment
-The notion of differential privacy (DP) Dwork et al.
-(2014) has been adopted in the alignment frameworks to ensure that the presence or absence of a single sample in a preference dataset does not significantly alter the outcome of the model.
+The notion of differential privacy (DP) dwork2014algorithmic has been adopted in the alignment frameworks to ensure that the presence or absence of a single sample in a preference dataset does not significantly alter the outcome of the model.
 Definition 1 ($(\epsilon,\delta)$ Differential Privacy).
@@
 However, when a labeler annotates multiple prompt-response pairs across the dataset, labeler-privacy guarantees become essential.
-This distinction between preference-privacy and labeler-privacy has been well recognized in the literature (McMahan et al., 2017; Liu et al., 2020; Levy et al., 2021).
+This distinction between preference-privacy and labeler-privacy has been well recognized in the literature (mcmahan2017communication; liu2020learning; levy2021learning).
 To extend preference-privacy guarantees to the labeler-privacy, privacy accounting and composition techniques can be adopted.
-For instance, if a labeler contributes to $k$ labeled examples in the dataset $\mathcal{D}$ with $(\epsilon,0)$-preference privacy, the Basic Composition theorem (Dwork et al., 2014) implies a labeler-privacy guarantee of $(k\epsilon,0)$.
+For instance, if a labeler contributes to $k$ labeled examples in the dataset $\mathcal{D}$ with $(\epsilon,0)$-preference privacy, the Basic Composition theorem (dwork2014algorithmic) implies a labeler-privacy guarantee of $(k\epsilon,0)$.
 To limit cumulative privacy loss, it is important to operate in high-privacy regimes (i.e., with small privacy budgets).
@@
 For instance, in dataset $\mathcal{D}$, if any labeler labels $k$ prompt-response pairs, with Basic Composition, $(\epsilon,0)$-Preference DP will satisfy $(k\epsilon,0)$-Labeler DP.
-However, more sophisticated methods, such as Advanced Composition (Dwork et al., 2014), Adaptive Composition Rogers et al.
-(2016), and the Moments Accountant Abadi et al.
-(2016) can be used to obtain tighter bounds depending on the application’s privacy requirements.
+However, more sophisticated methods, such as Advanced Composition (dwork2014algorithmic), Adaptive Composition rogers_privacy_2016, and the Moments Accountant abadi2016deep can be used to obtain tighter bounds depending on the application’s privacy requirements.
 With a small failure probability of $\delta^{\prime}$, Advanced composition provides ($\epsilon_{\text{Labeler}},\delta_{\text{Labeler}}$) Labeler DP where $\epsilon_{\text{Labeler}}=k_{i}\epsilon^{2}+\epsilon\sqrt{2k_{i}\log(\frac{1}{\delta^{\prime}})},$ and $\delta_{\text{Labeler}}=\delta^{\prime}$.
@@
 Though RR is simple to implement and ensures strong privacy guarantees, it introduces significant noise to the labels, which can degrade alignment quality, especially in small datasets or high-privacy regimes.
-Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al., 2016) ensures $(\epsilon,\delta)$-DP by gradient perturbation during alignment.
+Differentially Private Stochastic Gradient Descent (DP-SGD) (abadi2016deep) ensures $(\epsilon,\delta)$-DP by gradient perturbation during alignment.
 In each round, the gradients $(\mathbf{\bar{g}})$ are clipped with a clipping threshold $(C)$ and perturbed with Gaussian Noise $(\mathcal{N}(0,{\sigma}^{2}{C}^{2}I))$ where, noise scale $\sigma=\sqrt{2\log(1.25/\delta)}/\epsilon$.
@@
 Our key idea is to preserve the privacy of the labeler generated preferences during the multi-stage alignment process for better privacy-utility trade-off.
-To facilitate understanding, we first describe PROPS in a two-stage $(K=2)$ setting and the generalization to arbitrary number of stages is straightforward.
+To facilitate understanding, we first describe PROPS in a two-stage $(K=2)$ setting.
 We begin with the preference dataset $\mathcal{D}$ consists of $n$ samples.
@@
 Let us denote the corresponding label obtained from the model $M_{1}$ for a prompt as $\ell_{M_{1}}$.
-To summarize, at this point, for each tuple $(x,y_{1},y_{2})$, we have access to $(\ell_{RR},\ell_{M_{1}})$.
-Viewing these as two noisy observations of the ground truth $\ell^{*}$, it is natural to ask the following: can one combine these predictions to obtain a better prediction about the ground truth preference?
+To summarize, at this point, for each tuple $(x,y_{1},y_{2})$ in the partition $\mathcal{D}_{2}$, we have access to two noisy observations: its RR-perturbed label $\ell_{RR}$ (also from $\mathcal{D}_{2}$) and the model’s prediction $\ell_{M_{1}}$.
 Note that we know the error rate of RR ($\gamma_{\epsilon}$), however, we don’t know the error rate of the model $M_{1}$ (say $\gamma_{M_{1}}$, which denotes the probability with which the model $M_{1}$ makes errors).
@@
 In fact, it is not too difficult to work out the MLE combiner using two noisy observations.
+Note that these two noisy observations are fundamentally different.
+The $l_{RR}$ (Privacy Noise) is from the Randomized Response (RR) mechanism, and its flip probability, $\gamma_{\epsilon}$, is known, analytic, and fixed by the privacy budget $\epsilon$.
+In contrast, $l_{M_{1}}$ (Model generated label) is the prediction from the Stage-1 model; its error rate, $\gamma_{M_{1}}$, is unknown, empirical, and must be estimated, as it reflects the quality of the Stage-1 alignment.
 Assuming that the RR noise and the noise induced by the model $M_{1}$ are independent, the MLE statistic (log-likelihood ratio) can be written and simplified as follows:
-$\small\Lambda(\ell_{RR},\ell_{M_{1}})$ $=\log\left(\frac{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=0)}{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=1)}\right)$
-$=(-1)^{\ell_{RR}}\cdot\log\left(\frac{1-\gamma_{\epsilon}}{\gamma_{\epsilon}}\right)+(-1)^{\ell_{M_{1}}}\cdot\log\left(\frac{1-\gamma_{M_{1}}}{\gamma_{M_{1}}}\right).$ (4)
+$\Lambda(\ell_{RR},\ell_{M_{1}})=\log\left(\frac{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=0)}{\mathbb{P}(\ell_{RR},\ell_{M_{1}}\mid\ell^{*}=1)}\right)=(-1)^{\ell_{RR}}\cdot\log\left(\frac{1-\gamma_{\epsilon}}{\gamma_{\epsilon}}\right)+(-1)^{\ell_{M_{1}}}\cdot\log\left(\frac{1-\gamma_{M_{1}}}{\gamma_{M_{1}}}\right).$ (4)
 The above then yields the methodology one can use for creating a new label for each prompt as:
@@
 In addition, in Section A.3, we provide experimental evidence to verify the efficacy of our proposed approach and the assumption about $M_{1}$, by comparing our proposed estimation procedure with the scenario where an oracle provides access to the true labels.
-Our results indicate that our procedure is valid, as the gap between the estimated and “ground-truth” $\gamma_{M}$ is consistently small across various privacy budgets.
+Our results indicate that our procedure is valid, as the gap between the estimated and “ground-truth" $\gamma_{M}$ is consistently small across various privacy budgets.
 3.1 Theoretical Results for PROPS
 In this section, we analyze the Sub-optimality gap, which captures the gap between the optimal non-private DPO policy parameters $\theta^{*}$, and the policy parameters obtained through two-stage PROPS $\hat{\theta}_{\text{PROPS}}$.
-To obtain a bound on the Sub-optimality gap, we follow similar assumptions as those made in Chowdhury et al.
-(2024a), including expressing the model as a log-linear policy and assuming smoothness by placing bounds on the policy and its gradients.
+To obtain a bound on the Sub-optimality gap, we follow similar assumptions as those made in chowdhury2024provably, including expressing the model as a log-linear policy and assuming smoothness by placing bounds on the policy and its gradients.
 This gap is formalized in Theorem 1; we provide more details on the assumptions and derivation in Section A.5.
@@
 Algorithm 1 PROPS: PROgressively Private Self-alignment
-Inputs: Fine-tuned Model $M_{0}$, Dataset $\mathcal{D}$, Stages $K$, Privacy parameters $(\epsilon,\delta)$
-Output: Aligned model $M_{K}$
- Perform RR on $\mathcal{D}$, such that $\mathcal{D}\xrightarrow{RR(\gamma_{\epsilon})}\mathcal{D}^{\prime}$
- Partition the dataset $\mathcal{D}^{\prime}$ into $K$-disjoint datasets such that $\mathcal{D}^{\prime}=D_{1}\cup D_{2}\cup\ldots\cup D_{K}$
- Flip labels using Randomized Response (RR) with probability $\gamma_{\epsilon}=\frac{1}{1+e^{\epsilon}}$
- Align model $M_{0}$ with dataset $D_{1}$ to obtain $M_{1}$
-for $k=2,3,\ldots,K$ do
-  Generate labels $\ell_{M_{k-1}}^{D_{k}}$ for dataset $D_{k}$ using the model $M_{k-1}$ (from previous stage)
-  Obtain $\ell_{RR}^{D_{k}}$ and $\ell_{M_{k-1}}^{D_{k}}$ and obtain the maximum likelihood estimator (MLE) $\Lambda$ according to equation 9 and generate label as:
-$\small\ell_{\text{PROPS}}^{D_{k}}=\begin{cases}1,\quad\text{ If }\Lambda\leq 0\\ 0,\quad\text{ If }\Lambda>0.\end{cases}$
-  Align model $M_{k-1}$ on dataset $D_{k}$ with PROPS labels ($\ell_{\text{PROPS}}^{D_{k}}$) to get model $M_{k}$
-end for
-Output: Aligned model $M_{K}$ on dataset $\mathcal{D}$
+1:Fine-tuned model $M_{0}$, dataset $\mathcal{D}$, number of stages $K$, privacy parameters $(\epsilon,\delta)$
+2:Aligned model $M_{K}$
+3:Perform RR on $\mathcal{D}$ to obtain $\mathcal{D}^{\prime}$ $\triangleright$ $\mathcal{D}\xrightarrow{RR(\gamma_{\epsilon})}\mathcal{D}^{\prime}$
+4:Partition $\mathcal{D}^{\prime}$ into $K$ disjoint subsets: $\mathcal{D}^{\prime}=D_{1}\cup D_{2}\cup\dots\cup D_{K}$
+5:Set flip probability $\gamma_{\epsilon}=\dfrac{1}{1+e^{\epsilon}}$
+6:Align $M_{0}$ on $D_{1}$ (using $R$ labels) to obtain $M_{1}$
+7:for $k=2$ to $K$ do
+8:  Use $M_{k-1}$ to generate labels $\ell_{M_{k-1}}^{D_{k}}$ on $D_{k}$
+9:  Obtain $\ell_{\mathrm{RR}}^{D_{k}}$ and $\ell^{D_{k}}_{M_{k-1}}$ for dataset $D_{k}$ and obtain Maximum Likelihood Estimator (MLE) $\Lambda$ according to Equation 5, and generate labels as $\ell_{\text{PROPS}}^{D_{k}}=\begin{cases}1,&\text{if }\Lambda\leq 0,\\ 0,&\text{if }\Lambda>0,\end{cases}$
+10:  Align $M_{k-1}$ on $D_{k}$ with PROPS labels $\ell_{\text{PROPS}}^{D_{k}}$ to obtain $M_{k}$
+11:end for
+12:return $M_{K}$
 Remark 1 PROPS for RLHF.
@@
 We present the detailed adaption of PROPS for RLHF in Section A.5.
-Remark 2 Distinction from Label-DP, Multi-Stage RR and PATE: While the notion of ($\epsilon,\delta$)-Preference Privacy is motivated from the notion of $(\epsilon,\delta)$-Label DP (Chaudhuri & Hsu, 2011; Ghazi et al., 2021a), there are distinctions between the two frameworks.
+Remark 2 Distinction from Label-DP, Multi-Stage RR and PATE: While the notion of ($\epsilon,\delta$)-Preference Privacy is motivated from the notion of $(\epsilon,\delta)$-Label DP (chaudhuri2011sample; NEURIPS2021_e3a54649), there are distinctions between the two frameworks.
 In Label-DP, only labels are treated as private, and noise is added in proportion to the privacy budget to preserve the privacy of the dataset’s labels.
-In addition, RR (Warner, 1965) is a direct approach to implement preference privacy, as the preference $\ell_{i}$ of a data entry is flipped with probability $\gamma_{\epsilon}=\frac{1}{1+e^{\epsilon}}$.
-While PROPS shares similarities with Multi-Stage RR Ghazi et al.
-(2021b) and PATE Papernot et al.
-(2018), it differs significantly in approach and application.
+In addition, RR (warner1965randomized) is a direct approach to implement preference privacy, as the preference $\ell_{i}$ of a data entry is flipped with probability $\gamma_{\epsilon}=\frac{1}{1+e^{\epsilon}}$.
+While PROPS shares similarities with Multi-Stage RR ghazi2021deep and PATE papernot2018scalable, it differs significantly in approach and application.
 Unlike Multi-Stage RR, which relies on simple sampling for combining noisy labels and model predictions, PROPS uses an MLE-based approach for principled integration.
@@
 4 Experiments and Discussion
-Datasets and models: In our experiments and validation, we have used (1) three datasets (jondurbin/truthy-dpo-v0.1, Anthropic HH-RLHF, and AlpacaEval) and (2) three different models of varying sizes: Pythia-1B EleutherAI (2024) , GPT2-Large (774M) OpenAI (2024a) and GPT2-Medium (355M) OpenAI (2024b).
-We have adopted a similar experimental setup as the prior works Rafailov et al.
-(2024); Chakraborty et al.
-(2024); von Werra et al.
-(2020)Lior-Baruch (2024).
+Datasets and models: In our experiments and validation, we have used (1) three datasets (jondurbin/truthy-dpo-v0.1, Anthropic HH-RLHF, and AlpacaEval) and (2) three different models of varying sizes: Pythia-1B Pythia-1b , GPT2-Large (774M) GPT2-Large and GPT2-Medium (355M) GPT2-Medium.
+We have adopted a similar experimental setup as the prior works rafailov2024directpreferenceoptimizationlanguage; chakraborty2024transferqstarprincipled; vonwerra2022trllior-baruch.
 The code for PROPS is publicly available [^1] https://anonymous.4open.science/r/PROPS-2025.
@@
 PROPS vs DPSGD: In Table 2, we present Win-Tie rates comparing our proposed algorithm PROPS with the conventional DP-SGD algorithm for GPT2-Large and GPT2-Medium models on the truthy-dpo-v0.1 dataset.
-DP-SGD was ran using the Gaussian mechanism for 1 epoch with $\delta=10^{-10}$, a gradient clipping threshold of 10, and a batch size of 4 for 2 epochs.
+DP-SGD was ran using the Gaussian mechanism for 1 epoch.
+We set a gradient clipping threshold C = 10, as a lower value (e.g., C=1) introduced significant clipping bias that harmed utility.
+Our choice of $\delta=10^{-10}$ is intentionally conservative and stricter than the common $1/N$ heuristic to establish a strong privacy baseline.
 As the results indicate, PROPS is able to consistently outperform DP-SGD at higher privacy regimes ($\epsilon=0.1,0.5,1$) for both models.
@@
 $2.0$ $38.4$ $17.6$ $44.0$
+Results on multi-stage PROPS:
+We now report results comparing 2-stage and 3-stage PROPS using the truthy-dpo-v0.1 dataset with GPT2-Large in Table 3.
+Our findings highlight an interesting tradeoff in the PROPS pipeline: the optimal number of alignment stages depend on the available privacy budget.
+When the privacy budget is low (small $\epsilon$), each human-labeled preference needs to be heavily perturbed, resulting in high label noise.
+Under this regime, increasing the number of stages provides diminishing returns: early-stage models are themselves noisy, and relabeling through additional stages can further propagate this noise.
+Moreover, since the dataset must be partitioned across stages, later stages have less effective data to counteract the noise.
+As a result, a smaller number of stages (e.g., $K=2$) tends to yield better alignment in high-privacy settings.
+In contrast, when the privacy budget is moderate or higher (larger $\epsilon$), each preference label is less noisy, and earlier-stage models produce more reliable pseudo-labels.
+This improves the quality of relabeling in subsequent stages, allowing additional stages to refine alignment rather than amplify errors.
+Consequently, a larger number of stages (e.g., $K=3$) can begin to provide measurable gains, as evidenced in Table 3, where at $\epsilon=2.0$ the 3-stage setup achieves $44.0\%$ wins compared to $38.4\%$ for 2 stages.
+These results suggest that number of stages in PROPS should scale with privacy budget and data availability: fewer stages are more robust under strong privacy constraints and limited samples, whereas additional stages can be beneficial once the signal-to-noise ratio of the preference data improves.
 Figure 4: Prompt-Response pairs generated by PROPS and DP-SGD based GPT2-Large models and their corresponding scores (helpfulness and harmlessness).
@@
 More prompt-response examples are in Section A.7 of the appendix.
-Results on multi-stage PROPS: We now report results to observe if performance gains can be attained by increasing the number of stages used in PROPS.
-In Table 3, we present results comparing 2-stage and 3-stage PROPS using the truthy-dpo-v0.1 dataset with GPT2-Large.
-Results indicate that across high privacy budgets, 2-stage PROPS outperforms 3-stage PROPS.
-This implies that further hyperparameter tuning under privacy constraints is required for 3 or more stages.
-We leave this as future work as our results up to this point have indicated that the 2-stage results effectively illustrate the key performance trends under varying privacy constraints.
 Illustrative Example Responses to Prompts for varying privacy levels: To illustrate the privacy-utility tradeoff, responses from LLMs trained with different privacy levels ($\epsilon=0.1,1,\infty$) were compared as shown in Figure 4.
@@
 By improving the privacy-utility tradeoff, this approach supports the use of such language models while fostering trust in systems that rely on human data.
-Due to resource constraints, we have used “smaller” language models (e.g.
+Due to resource constraints, we have used “smaller" language models (e.g.
 GPT2-Medium, GPT2-Large, Pythia-1B); however, our results still indicate the effectiveness of PROPS in ensuring preference level privacy.
@@
 This work contributes positively to the field by promoting ethical considerations in LLM training and alignment.
+5.1 Acknowledgments
+This work was supported by NIH Award R01-CA261457-01A1 and also by the US Department of Energy, Office of Science, Office of Advanced Scientific Computing under Award Number DE-SC-ERKJ422, and US NSF under Grants CCF-2100013, CNS-2209951, CNS-2317192.
 Appendix A Appendix
@@
 This 75% segment was divided into two halves, with three epochs of DPO run on each half.
-Subsequently, the dataset was filtered to include only preference pairs generated by prompting a Large Language Model (LLM) to act as ”an honest and helpful assistant.” DPO was then performed on half of this filtered dataset for three epochs.
+Subsequently, the dataset was filtered to include only preference pairs generated by prompting a Large Language Model (LLM) to act as "an honest and helpful assistant." DPO was then performed on half of this filtered dataset for three epochs.
 Win-Tie-Loss rates were calculated using the remaining 10% of the Truthy-DPO-v0.1 dataset, which consists of 100 prompts.
@@
 PROPS incurs the highest computational cost, as it integrates several components: the baseline DPO cost, the randomized response mechanism, model inference over a second data partition, and an additional phase of maximum likelihood estimation (MLE).
+Comparison of Wall-Clock Time : While Big-O analysis provides theoretical insights into computational complexity, empirical comparisons of wall-clock time are crucial for understanding the practical overhead introduced by PROPS.
+We benchmarked a standard DPO run against our two-stage PROPS procedure using GPT-2 Medium and GPT-2 Large models, with results summarized in Table 5.
+For a given model and dataset size, the core DPO training times are nearly identical across methods.
+The primary additional cost in PROPS arises from Stage 2, where pseudo labels ($l_{M_{1}}$) are generated on the second data partition ($D_{2}$).
+In our experiments, this pseudo-label generation constitutes the dominant component of the overhead.
+Nevertheless, this extra computation represents a necessary and well-justified trade-off.
+As shown by the win-rate results in Section 4 (e.g., Tables 1) and Appendix A.6 (Tables 7), the inclusion of this step enables PROPS to achieve markedly higher model utility and alignment quality compared to baselines such as RR and DP-SGD—particularly under stringent privacy constraints ($\epsilon\leq 1$) where these alternatives exhibit substantial performance degradation.
+In essence, the additional compute directly translates into meaningful improvements in model performance and alignment.
+Table 5: Wall-clock time comparison (500 entries)
+Model Method DPO Training Label Generation Total Time
+GPT2-M $R$ 42 s 0 s 42 s
+GPT2-M PROPS 42 s 4 min 31 s 5 min 13 s
+GPT2-L $R$ 1 min 53 s 0 s 1 min 53 s
+GPT2-L PROPS 1 min 53 s 11 min 36 s 13 min 29 s
+While PROPS incurs additional computational cost due to the label generation step, this overhead is justified by its substantially improved alignment quality, particularly in high-privacy settings.
+As demonstrated in our main experimental results, PROPS achieves significantly higher win rates compared to RR and DP-SGD under stringent privacy budgets ($\epsilon\leq 1$), indicating that the extra computation translates directly into better model utility where conventional methods falter.
 A.2 MLE estimator for $\ell^{*}$ using $(\ell_{RR},\ell_{M_{1}})$
@@
 The use of disjoint data across stages reduces the risk of direct dependence between the two datasets.
-Validation of our estimation procedure for $\gamma^{*}_{M_{1}}$: To validate the robustness of our estimation process for obtaining $\hat{\gamma_{M_{1}}}$, we compare the estimated error rate of model $M_{1}$ with its true counterpart (the ”oracle” error rate of $\gamma_{M_{1}}^{*}$).
+Validation of our estimation procedure for $\gamma^{*}_{M_{1}}$: To validate the robustness of our estimation process for obtaining $\hat{\gamma_{M_{1}}}$, we compare the estimated error rate of model $M_{1}$ with its true counterpart (the "oracle" error rate of $\gamma_{M_{1}}^{*}$).
 Specifically, Table 6 presents a comparison between the estimated error rate $\hat{\gamma}_{M_{1}}$, computed using equation 8 from the paper, and the oracle error rate $\hat{\gamma}_{M_{1}}^{*}$, obtained by evaluating $M_{1}$’s predictions on the unperturbed preference data against the ground truth (i.e., the original preferences prior to flipping).
 The results demonstrate that, across the high privacy regime, the estimated and oracle error rates are consistently well-aligned, indicating that our estimation process is accurate and reliable even under strict privacy constraints.
-Table 5: Comparable estimations on flipping for our method ($\hat{\gamma_{M_{1}}}$) compared to the model flipping probability obtained from the ”oracle” setting $\gamma_{M_{1}}^{*}$.
+Table 6: Comparable estimations on flipping for our method ($\hat{\gamma_{M_{1}}}$) compared to the model flipping probability obtained from the "oracle" setting $\gamma_{M_{1}}^{*}$.
 Privacy Budget $(\epsilon)$ $\hat{\gamma_{M_{1}}}$ (Our Method) $\gamma^{*}_{M_{1}}$ (Obtained from Oracle)
@@
 A.4 Proof Sketch of Theorem 1: Sub-optimality gap for PROPS
-Chowdhury et al.
-(2024a) provides a bound on the sub-optimality gap between the rewards obtained using an optimal model aligned under DPO with noiseless preference data and a model aligned under their proposed robust DPO (rDPO) method, which accounts for noisy (i.e.
+chowdhury2024provably provides a bound on the sub-optimality gap between the rewards obtained using an optimal model aligned under DPO with noiseless preference data and a model aligned under their proposed robust DPO (rDPO) method, which accounts for noisy (i.e.
 flipped) preference data.
 Their result characterizes how many preference samples are needed at different noise levels to ensure that the loss in rewards (relative to the original DPO method) does not exceed a certain bound.
-Chowdhury et al.
-(2024a) assume that the model can be expressed as a log-linear policy (i.e.
+chowdhury2024provably assume that the model can be expressed as a log-linear policy (i.e.
 as a function of a feature map $\phi(x,y)^{T}$ and the weights of the last layer $\theta$).
@@
 Therefore, the flipping rate of PROPS can be described as $\gamma_{\text{MLE}}=\min(\gamma_{\epsilon},\gamma_{M_{1}})$.
-Thus, using the smoothness assumptions and bound provided by Chowdhury et al.
-(2024a), the sub-optimality gap between between the optimal non-private DPO policy parameters $\theta^{*}$ and the policy parameters obtained through two-stage PROPS $\hat{\theta}_{\text{PROPS}}$ can be obtained as:
+Thus, using the smoothness assumptions and bound provided by chowdhury2024provably, the sub-optimality gap between between the optimal non-private DPO policy parameters $\theta^{*}$ and the policy parameters obtained through two-stage PROPS $\hat{\theta}_{\text{PROPS}}$ can be obtained as:
 $\small\underbrace{\left\lVert\hat{\theta}_{\text{PROPS}}-\theta^{*}\right\rVert}_{\text{Sub-Optimality Gap}}\leq\mathcal{O}\left(\frac{\sqrt{\kappa}}{\gamma\beta(1-2\cdot\min(\gamma_{M_{1}},\gamma_{\epsilon}))}\sqrt{\frac{d}{n_{2}}}\right),$
 where $n_{2}$ is the number of samples used in stage 2 of PROPS.
+Limitations of Theoretical Analysis: Our theoretical analysis relies on idealized assumptions—such as log-linear policies, smoothness, and feature coverage—that enable tractable mathematical treatment but may not hold for modern Transformer-based architectures.
+These simplifying assumptions are standard in the literature and are intended to provide interpretable intuition about the underlying mechanisms rather than exact quantitative guarantees.
+In particular, the theory correctly predicts how performance scales with key parameters such as the effective noise rate of the first-stage MLE ($\min(\gamma_{M_{1}},\gamma_{\epsilon})$) and the second-stage sample size ($n_{2}$), offering a principled explanation of the observed trade-offs.
+As a result, while the derived scaling trends offer valuable insight and a principled sanity check on empirical behavior, the bounds should not be interpreted as directly predictive of large-scale model performance.
+Bridging this gap between theoretical idealizations and real-world LLM architectures remains an important direction for future work.
 A.5 Adapting PROPS for RLHF-based Alignment
@@
 RR and PROPS vs.
-DP-SGD in Tables 6 and 7 respectively.
+DP-SGD in Tables 7 and 8 respectively.
 Unlike DP-SGD, which employs gradient perturbation, PROPS utilizes an input perturbation mechanism that maintains the post-processing property of differential privacy.
@@
 Figure 6 shows an example of how these measurements were obtained.
-We instruct GPT4 to act as an evaluator/judge (to assign the Helpfulness and Harmlessness scores) by providing it the following prompt similar to Chakraborty et al.
-(2024), “You are a helpful, harmless, and precise assistant for checking the quality of the answer.
+We instruct GPT4 to act as an evaluator/judge (to assign the Helpfulness and Harmlessness scores) by providing it the following prompt similar to chakraborty2024transferqstarprincipled, “You are a helpful, harmless, and precise assistant for checking the quality of the answer.
 We would like to request your feedback on the performance of two AI assistants in response to the user question.
@@
 The two scores are separated by a space.
-In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.” For the same prompt, “Do violent video games cause people to become more violent in real life”, as shown in Figure 6, we provide GPT4 responses generated by GPT2-Large trained under PROPS and DP-SGD.
+In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment." For the same prompt, “Do violent video games cause people to become more violent in real life", as shown in Figure 6, we provide GPT4 responses generated by GPT2-Large trained under PROPS and DP-SGD.
 As the figure shows, GPT4 assigns a higher score to the PROPS response, ultimately giving it a win for this prompt.
@@
 Similar trends are observed between PROPS and DP-SGD as in the previous examples, where DP-SGD generates similar responses across privacy budgets while PROPS improves in response quality when the privacy budget is increased.
-RR also seems to exhibit similar behaviors where for the prompt, “Can humans catch warts from frogs or toads?”, it generates a more focused response at $\epsilon=0.1$ rather than $\epsilon=1$.
+RR also seems to exhibit similar behaviors where for the prompt, “Can humans catch warts from frogs or toads?", it generates a more focused response at $\epsilon=0.1$ rather than $\epsilon=1$.
 However, looking at $\epsilon=0.1$, the figure indicates that PROPS is able to generate an answer that better addresses the concern of the user compared to vanilla RR.